### v1

# NeuroSim V1.5: Improved Software Backbone for Benchmarking Compute-in-Memory Accelerators with Device and Circuit-level Non-idealities

Abstract
The exponential growth of artificial intelligence (AI) applications has exposed the inefficiency of conventional von Neumann architectures, where frequent data transfers between compute units and memory create significant energy and latency bottlenecks. Analog Computing-in-Memory (ACIM) addresses this challenge by performing multiply-accumulate (MAC) operations directly in the memory arrays, substantially reducing data movement. However, designing robust ACIM accelerators requires accurate modeling of device- and circuit-level non-idealities. In this work, we present NeuroSim V1.5, introducing several key advances: (1) seamless integration with TensorRT’s post-training quantization flow enabling support for more neural networks including transformers, (2) a flexible noise injection methodology built on pre-characterized statistical models, making it straightforward to incorporate data from SPICE simulations or silicon measurements, (3) expanded device support including emerging non-volatile capacitive memories, and (4) up to 6.5× faster runtime than NeuroSim V1.4 through optimized behavioral simulation. The combination of these capabilities uniquely enables systematic design space exploration across both accuracy and hardware efficiency metrics. Through multiple case studies, we demonstrate optimization of critical design parameters while maintaining network accuracy. By bridging high-fidelity noise modeling with efficient simulation, NeuroSim V1.5 advances the design and validation of next-generation ACIM accelerators. All NeuroSim versions are available open-source at https://github.com/neurosim/NeuroSim.
Index Terms:
Compute-in-memory, AI accelerator, hardware/software co-design, open-source, benchmark
I Introduction
The exponential growth in AI applications has exposed a critical challenge in energy-efficient computing. The fundamental limitation lies in the mismatch between the memory-centric workloads such as in machine learning algorithms and the processing-centric architecture in contemporary hardware platforms. This mismatch causes massive data movement between memory and processing units to consume more than five times the energy of computation [1].
Figure 1: Overview of NeuroSim V1.5: This update revamped the software backbone to support arbitrary neural network models and provide a more intuitive noise simulation workflow tailored to device and circuit experts. NeuroSim V1.5 enables users to evaluate inference accuracy across diverse architectures, including custom convolutional neural networks and transformers, while simulating various device types and noise sources. The platform uniquely supports power, performance, and area (PPA) estimations, facilitating comprehensive system-level design space exploration. Key new features include optimized noise modeling, compatibility with advanced non-volatile memory technologies, and improved runtime performance.
The challenges stem from the data-intensive nature of deep neural networks (DNNs), where a relatively small number of operations is performed per fetched data; this causes system performance to be limited by memory bandwidth, as noted by the roofline model [2]. Compute-in-memory (CIM) has emerged as a promising paradigm to tackle this bottleneck by co-locating data storage and multiply-accumulate (MAC) operations in the same physical arrays, thereby reducing repeated weight fetches. In digital computing-in-memory (DCIM), memory arrays (typically made of modified SRAM bit cells) are augmented with digital multipliers and adder trees for partial-sum accumulation, delivering robust performance with the overhead of adder trees [3].
By contrast, ACIM leverages the intrinsic electrical behavior of memory cells to carry out multiplications directly in the current or charge domain. Specifically, weight values are encoded as conductance (or capacitance), and when an input voltage is applied to a row, the resultant column current (or charge) naturally represents the multiplication of input and weight. Summing all such column outputs yields the MAC result in the analog domain and then being converted back to the digital domain through analog-to-digital converters (ADCs) [4, 5]. A variety of emerging non-volatile memory (NVM) devices have been utilized for these ACIM arrays, including resistive random-access memory (RRAM) [6, 7, 8, 9], phase-change memory (PCM) [10, 11, 12], ferroelectric field-effect transistors (FeFET) [13, 14, 15], and more recently non-volatile capacitors (nvCap) [16, 17, 18]. ACIM implementations can also rely on modified SRAM bit cells, tapping into current-based summation [19, 20, 21, 22, 23, 24], or incorporate on-chip capacitors to accumulate charge [25, 26] or using them for time-domain computing [27]. Several of these memory cells are depicted in a generic CIM macro with peripheral circuits as shown in Fig. 1.
Over the past few years, both DCIM and ACIM have been demonstrated in silicon with memory capacities up to a few megabytes [28]—sufficient for many on-chip inference scenarios such as computer vision tasks involving millions of parameters. DCIM generally offers the robustness of digital processing, making it attractive for high-performance computing, whereas ACIM excels in energy efficiency, an essential requirement for power-constrained edge devices.
ACIM accelerators can theoretically surpass 100 TOPS/W [3, 9, 19, 26]—far exceeding the 1–10 TOPS/W range of contemporary von Neumann platforms—largely because they eliminate most weight-reloading costs by performing the MAC operations directly within the memory arrays. In this paper, one operation is defined as an 8b×8b MAC. The key challenge in ACIM is to manage inherently noisy analog computations to maintain the inference accuracy.
These noise sources span multiple abstraction layers. At the device level, fabrication imperfections lead to device-to-device (D2D) variations and stuck-at-faults. Memory states can also drift over time, degrading compute accuracy [29, 30, 31, 32]. At the circuit level, thermal noise and process variations in peripheral circuits degrade sensing margins, resulting in errors in the ADC processes [33, 18]. At the system level, partial sums may be quantized at a low bit precision, compounding errors as layer outputs are passed from one stage to another [34, 35, 36]. Together, these effects can reduce the inference accuracy of DNNs run on ACIM chips. Consequently, modeling and ultimately mitigating these non-idealities is critical for the design of robust ACIM architectures.
These multi-layered noise effects create complex interactions between device characteristics, circuit design choices, and neural network requirements. Therefore optimizing ACIM accelerators requires systematic exploration across this design space to balance inference accuracy with hardware efficiency. Such an exploration requires simulation frameworks that can evaluate both accuracy and hardware metrics while maintaining practical runtimes even for modern large-scale networks.
Direct SPICE simulation of analog computing arrays is prohibitively expensive for large neural networks, making it impractical to evaluate broad parameter spaces with modern datasets. While tools like CrossSim [37] and AIHWKit [38] offer detailed device modeling and noise-aware training respectively, they lack integrated hardware analysis capabilities and require extensive engineering effort to support new devices or circuits. Other frameworks like CIMLoop [39] focus on hardware metrics without considering accuracy under device and circuit non-idealities. These limitations make it difficult to systematically explore and optimize ACIM designs across both software accuracy and hardware efficiency metrics.
NeuroSim [40] addresses these challenges through an integrated framework that combines behavioral simulation of CIM operation with automated hardware analysis. Building on this foundation, we present NeuroSim V1.5, which introduces several key advances to enable comprehensive design space exploration:
  1. 1.
    Seamless integration with TensorRT to support arbitrary pre-trained models, while automating quantization.
  2. 2.
    A flexible noise modeling methodology built on pre-characterized statistical models, enabling rapid evaluation of diverse device and circuit designs.
  3. 3.
    Support for emerging technologies like charge-based non-volatile capacitive memories (nvCap) and emerging algorithms like vision transformer (ViT).
  4. 4.
    Significantly improved runtime performance - up to 6.5× faster than previous versions of NeuroSim.
These capabilities maintain seamless compatibility with NeuroSim’s existing power, performance, and area (PPA) analysis framework, creating a unique environment for systematic CIM accelerator design optimization. The framework achieves runtime only two to three times that of software-only standard GPU inference.
The remainder of this paper is organized as follows: Section II provides technical background on ACIM systems and existing simulation approaches. Section III details our improved framework architecture and generic noise modeling methodology. Section IV presents results from various case studies including inference accuracy analysis of several devices and compute circuits using data from SPICE simulations and tapeout measurements [41, 42, 43, 33, 18, 7]. We also provide design space exploration for a RRAM based ACIM chip to demonstrate the framework’s adaptability and improved running speed. Finally, Section VI discusses implications for future larger-scale AI model benchmarking and concludes the paper.
II Background Information
II-A CIM Array Architecture and Operation
ACIM performs matrix computations by utilizing physical properties of memory arrays. The fundamental building block is the crossbar array, where memory elements are arranged at the intersections of wordlines (WLs) and bitlines (BLs).
This crossbar arrangement is uniquely suited for parallel MAC operations because each row’s input signal (voltage) is simultaneously applied to all memory devices along that row. The resulting outputs (currents or charges) are summed along each column, naturally implementing the dot product. Consequently, large-scale MAC operations can be performed in one analog ‘step’, vastly increasing the throughput compared to digital architectures that must shuttle data back and forth between separate memory and logic units. To date, two primary methods of analog computation have emerged.
Resistive arrays constructed with RRAM, PCM, FeFET that encode neural network weights as conductance states (G), performing MAC operations in the current domain:
$I_{BL_{i}}=\sum_{j}G_{ij}V_{j}$ (1)
where $I_{BL_{i}}$ represents the analog matrix-vector multiplication (MVM) output current on the $i^{th}$ bitline, and $j$ indexes across the wordlines.
Capacitive arrays utilize non-volatile capacitors (nvCap), where the weight is represented by the programmable capacitance of the ferroelectric or charge-trap memory cell [44] [45], thus performing MAC operations in the charge domain:
$Q_{BL_{i}}=\sum_{j}C_{ij}V_{j}$ (2)
Alternative charge-based computation approaches utilize fixed capacitors in conjunction with SRAM cells, where the stored digital weight modulates charge accumulation through charge sharing mechanisms [46, 47].
DCIM has emerged as another promising approach [3], where SRAM arrays are augmented with digital multipliers and adder trees to perform integer MAC operations directly within memory. These systems are attractive due to their compatibility with advanced CMOS technology scaling and robust digital operation. However, the digital processing overhead - particularly from the complex adder trees required for partial sum accumulation - impacts array density. Additionally, SRAM leakage power becomes more significant in advanced technology nodes. Both the array scalability and leakage power create challenges for running complete neural networks on-chip.
Each approach is finding its niche, while ACIM remains near the theoretical limit of efficiency for neural network applications. In ACIM designs, the choice of memory technology influences key array characteristics like precision, reliability, and energy efficiency - topics we will explore in detail when discussing noise modeling methodology.
II-B Device and Circuit Level Considerations
The precision and reliability of analog computing operations depend heavily on the properties of individual memory cells. Key device parameters include:
On/Off Ratio: The ratio between maximum and minimum conductance/capacitance states defines the usable dynamic range for weight storage. Finite on/off ratios introduce non-zero current/charge contribution from cells programmed to the off states, which accumulates across array columns and can degrade computation accuracy. This effect becomes particularly significant in larger arrays where hundreds of devices in the off-state contribute to the output. To address this, array designs typically include additional ‘dummy’ columns programmed to off states and are supported in NeuroSim [40], whose outputs are subtracted from the computational columns to cancel the accumulated off-state current/charge. While this technique effectively eliminates the impact of finite on/off ratio, it slightly reduces the available dynamic range for computation since part of the range must be reserved for cancellation. This results in smaller sensing margins and potentially smaller signal-to-noise ratio (SNR). Therefore, high on/off ratios are still a desired characteristic for ACIM memory devices.
Number of States: Supporting high-precision weights requires either multi-level memory cells (MLC) or bit slicing with multiple lower-precision/binary cells. For MLC devices, the number of reliably programmable states determines the weight precision. RRAM, PCM, and nvCap typically achieve 1-4 bits per cell [8, 28, 11], while FeFET have demonstrated >5 bits [14, 48].
At the circuit level, array size is limited mainly by sensing margins, where reliable output detection requires sufficient separation between computational states. This becomes challenging with larger arrays due to increased noise accumulation and reduced signal levels from parasitic effects.
Critical peripheral circuits include:
  • WL/BL drivers to supply the input and output signals.
  • ADCs for digitizing array outputs.
  • Shift-and-add circuits for managing bit-precise computations.
  • Accumulation units on the PE level for partial sum aggregation.
  • Multiplexers for data routing and array access.
The digital circuit blocks in the periphery can be verified through conventional VLSI design methodologies. Therefore, behavioral modeling efforts focus primarily on the analog computational elements.
Figure 2: Procedure for ACIM behavioral simulation in NeuroSim V1.5. (1) Inputs and weights are quantized from floating point to integer. (2) Inputs and weights unfolded and (3) mapped to compute-in-memory arrays. (4) Analog MACs are calculated with analytical circuit models. (5) Noise is injected either to memory devices or to digital MAC outputs. (6) ADC outputs are shifted and added. (7) Outputs from memory arrays are accumulated and concatenated. (8) Accumulated outputs are folded into output tensors. (9) Outputs are de-quantized from integer back to floating-point.
II-C Precision and Data Representation
Neural network parameters must be carefully mapped to CIM arrays while managing precision and computational accuracy. This mapping presents two key challenges: representing multi-bit input activations and storing high-precision weights. For input activations, two main approaches exist:
  1. 1.
    Using digital-to-analog converters (DACs) to represent each input value as an analog voltage. While straightforward, this requires complex multi-bit DACs for every row of the array, introducing significant area and energy overhead. The I-V or C-V nonlinearity of the underlying memory devices may also introduces the errors in the input representation.
  2. 2.
    Processing each bit of the digital input serially, with outputs shifted and added according to bit significance. This eliminates DAC overhead but requires multiple computational cycles.
For weight storage, two methods are available:
  1. 1.
    Bit slicing across multiple lower-precision cells. For example, an 8-bit weight can be distributed across cells of various precisions - from a single 8-bit cell to eight 1-bit cells, with intermediate combinations possible.
When slicing the weight among multiple cells, each cell is placed in an adjacent column, with the outputs of each column scaled by its significance post-ADC and added with the other scaled outputs. This operation is referred to as ‘shift-add’ and is described in Fig. 2 and in section III. Using analog shift-add before the ADCs has also been explored [44, 49, 50], but is not explicitly supported in NeuroSim V1.5.
While NeuroSim’s behavioral simulator supports both input representation methods, the PPA estimator only supports the bit-serial approach due to its practical hardware implementation.
In CIM, the fundamental MAC operation is performed in the integer domain, necessitating quantization of the inputs and weights of each layer in the network from floating-point to integer. The analog multiplication and accumulation processes generate integer results, which are digitized through ADCs. Finally, the integer outputs are de-quantized back to floating-point for further processing. Some accelerators explore quantizing the entire network to integers[51, 52], requiring specialized approximations for operations like normalization and activation functions. This approach is particularly applicable to transformers, which require more complex operators like GELU and layer normalization.
II-D Hierarchical System Architecture
The organization of ACIM accelerators stems from fundamental constraints in both device technology and neural network structure. Memory array sizes are typically limited to dimensions between 32×32 and 256×256 cells due to sensing margins and manufacturing considerations. However, neural network layers often require much larger matrix operations. This dimensional mismatch necessitates a hierarchical decomposition of network operations.
The organization of the chip hierarchy in NeuroSim is as follows: At the base level, synaptic arrays perform MAC operations through crossbar structures. These arrays are grouped into processing elements (PEs), which implement weight-stationary, spatial mapping of convolutional operations. For example, when mapping 3D convolutional kernels to 2D array structures, multiple arrays within a PE process different kernel depths in parallel, with their outputs appropriately accumulated.
PEs are further aggregated into computational tiles, each capable of processing a portion of a neural network layer. The tile architecture incorporates buffers for activation storage and accumulation circuits for combining partial results from multiple PEs. This organization enables efficient processing of large layers by distributing computations across multiple tiles while maintaining data locality. At the chip level, multiple tiles operate together through either an H-tree or X-Y bus interconnect structure. This highest level of hierarchy includes global buffers for inter-tile data movement and accumulation circuits for combining results from distributed computations. The entire system implements layer-wise pipelining, where different tiles process consecutive layers of the network simultaneously to maximize throughput.
To balance processing speed between layers of different sizes, weight duplication can be employed within tiles [53]. For instance, if a layer’s weight matrix is smaller than the tile’s capacity, the weights can be duplicated to process multiple input vectors simultaneously, improving throughput.
This hierarchical structure enables flexible mapping of diverse neural network architectures while managing the physical constraints of ACIM arrays. To model this hierarchy, NeuroSim employs analytical models for circuit blocks derived from standard cell libraries in advanced logic nodes. The previously released V1.4 [54] expanded technology support to the leading-edge 1 nm node, introducing stacked nanosheet transistors and studying the scaling path toward a future A5 node with complementary-FET (CFET) technology. Thanks to NeuroSim’s software/hardware co-design philosophy and open-source development model, the tool enables broader research on emerging device technologies, array architectures, on-chip training [55], and advanced packaging [56, 57, 58]. For a detailed analysis of the hierarchy in NeuroSim, readers are referred to [40].
III Simulation Methodology
III-A Overview of Design Methodology
Designing ACIM accelerators requires optimization across multiple abstraction layers - from device characteristics to system architectures. NeuroSim V1.5 provides an integrated framework to evaluate these designs through two main components shown in Fig. 1 working jointly:
  1. 1.
    A behavioral simulator that maps and quantizes neural networks for ACIM implementation and evaluates inference accuracy while accounting for hardware non-idealities.
  2. 2.
    A hardware analyzer that estimates system-level energy, latency, and area through detailed circuit models of array peripherals and interconnects.
In this update, our first contribution is a comprehensive re-design of the behavioral simulator. Second, we update the hardware analyzer to support the emerging nvCap-based CIM, complete with circuit models for the new charge-domain compute mode. In the behavioral simulator, we have fundamentally enhanced how we model compute-in-memory operations for accuracy estimation. A key approach is that we leverage pre-characterized statistical models derived from SPICE simulations or silicon measurements to ensure accuracy, while utilizing GPU-accelerated PyTorch operations to achieve the throughput needed for large-scale inference simulation. This separation allows us to maintain high fidelity while dramatically improving simulation efficiency.
The behavioral simulator supports two primary simulation modes:
Device expert mode: Model memory array physics directly, simulating conductance variations, stuck faults, and temporal drift to isolate the effect of memory technologies on network accuracy.
Circuit expert mode: Model memory arrays with a statistical approach using pre-characterized noise distributions from SPICE simulations or silicon measurements, enabling rapid evaluation of aggregated non-idealities from devices and circuits.
Through integration with TensorRT’s post-training quantization flow, NeuroSim V1.5 automatically maps pre-trained neural networks to the corresponding ACIM architecture.
The following sections detail the quantization methodology, array modeling approach, and specific techniques for incorporating device and circuit-level non-idealities into the simulation framework.
III-B Neural Network Quantization and Mapping
III-B1 Quantization
NeuroSim V1.5 adopts NVIDIA’s TensorRT quantization infrastructure for converting neural networks from floating-point to integer representations. This differs from previous NeuroSim versions which used in-house WAGE quantization [59] - an integer-training method that requires retraining the network. TensorRT instead provides post-training quantization through automatic operator replacement, supporting quantization of pre-trained networks using operators like linear layers, convolutions, and pooling without modification. TensorRT supports configurable bit widths for both inputs and weights, which we use to evaluate different precision settings (e.g., 8b/8b, 6b/6b).
The framework determines optimal scaling factors through either max or histogram calibration during a brief calibration phase. This phase typically requires only 2 batches of calibration data (e.g., several hundred images) to analyze the distribution of values in both input tensors and weight matrices. Max calibration simply uses the maximum observed value across these tensors to set the scaling factor. Histogram calibration creates a histogram of the observed tensor values and sets the scale factor based on the cumulative distribution function (CDF) - specifically, choosing the value at which the CDF reaches a user-specified percentile (typically 99.99%). This helps exclude outliers that would otherwise force a larger dynamic range. The key advantages of adopting TensorRT over WAGE quantization are direct support for arbitrary pre-trained models without retraining and the automated support for multiple operator types through operator replacement.
III-B2 Network to Array Mapping
Physical constraints of ACIM arrays necessitate careful partitioning of neural network operations. For a linear layer with input dimension $N$ and output dimension $M$, the quantized weight matrix must be mapped to multiple arrays based on both the memory array dimensions and the precision requirements. The number of memory cells required per weight depends on the bit precision of the cells ($b_{cell}$). For example, to represent an 8-bit weight using 1-bit cells, we need $N_{cell}=\lceil 8/1\rceil=8$ cells per weight. This bit-slicing requirement effectively multiplies the width dimension of our mapping by $N_{cell}$. As shown for a linear layer in Fig. 3, these weight components are arranged in adjacent columns, with each column corresponding to a different bit significance.
Memory array size further constrains the mapping. With a memory array height of $R$ rows and $C$ columns, the $N\times M$ weight matrix must be partitioned into $\lceil N/R\rceil$ rows of arrays. The total number of arrays needed to map the weights is equal to $\lceil N/R\rceil\cdot\lceil MN_{cell}/C\rceil$. Mapping convolutional layers to memory arrays requires additional unfolding operations, of which multiple methods can be implemented. For a more detailed description, readers are referred to [40].
The input feature map is unfolded and mapped to the arrays according to the selected input scheme. In Fig. 2 we show the bit-serial operation, where the feature map is processed separately for each bit of the quantized input.
Outputs from vertically-aligned arrays are accumulated, as these represent partial sums from the same output neurons. Outputs from horizontally aligned arrays are shifted and added according to their bit significance, then concatenated to form the complete layer output.
III-B3 Support for Transformers
For transformer architectures, NeuroSim V1.5 adopts a hybrid approach that maps compatible operations to CIM arrays while implementing other operations in the digital domain. The framework maps linear layers (including query, key, value projections and MLP layers) to CIM arrays using the same methodology as standard neural networks. These linear layers comprise a significant portion of the total computation in transformer models. However, dynamic operations such as attention score computation, layer normalization, and activation functions must still be computed.
We elect to compute these dynamic operations in floating-point compute units due to CIM hardware constraints. Non-volatile memory-based CIM is optimal for weight-stationary computations that do not require frequent weight updates. This is due to limited write endurance and slow programming speed common in these devices. Therefore, operators like self-attention that require unique input and weight matrices each inference are not well-suited for non-volatile CIM. Digital CIM implementations using SRAM could potentially handle these dynamic operations, however they would still require careful quantization of the query, key, and value matrices to maintain compatibility with integer-only computation [60]. Recent work on integer-only transformers [51, 52] has demonstrated promising approaches for full integer quantization, though further research is needed to evaluate the accuracy-efficiency tradeoffs of the hardware.
The current hardware estimation portion of NeuroSim does not include circuit models for floating-point functional units needed for self-attention computation, layer normalization, or GELU activation functions. Future versions will explore integrating these components, potentially through integer approximations that maintain accuracy while enabling efficient hardware implementation.
Figure 3: NeuroSim V1.5 flowchart. Device expert mode: user provides device data (memory state variation, stuck-at-faults, state drift). Circuit expert mode: user provides statistical data on MAC outputs from SPICE Monte-Carlo simulations or tapeout measurements.
III-C ACIM Behavioral Modeling
III-C1 Memory Array Operation
Building on TensorRT’s operator replacement strategy, NeuroSim V1.5 substitutes standard neural network layers with specialized CIM operators that model array-level computation. When a PyTorch model is loaded, linear and convolutional layers are automatically replaced with cim.Linear and cim.Conv modules that partition operations according to array constraints and simulate analog computation as described in Fig. 3.
The framework incorporates two separate simulation modes for modeling non-idealities: device expert mode for measured device-level device variation or calibrated aging models, and circuit expert mode using statistical models derived from SPICE or silicon measurements. These modes cannot be used simultaneously as circuit-level noise inherently captures device-level variations.
III-C2 Device and Circuit Non-Idealities
In device expert mode, integer weights are first mapped to conductance or capacitance states based on the target memory technology. The framework then simulates analog multiplication through Ohm’s law (I = GV) for resistive arrays or charge-based computation (Q = CV) for capacitive arrays. Input values are treated as idealized voltages, and outputs are computed by summing the resulting currents or charges along each column.
To capture all possible input and weight bit-slicing combinations, including both bit-serial and DAC-based input schemes, we use the following generalized equation:
$y=\sum_{i}^{N_{cell}}\sum_{j}^{N_{in}}(2^{i\cdot b_{cell}})(2^{j\cdot P_{DAC}})(W_{i}\cdot x_{j})$ (3)
where:
  • $b_{in}$ is the bit significance of the inputs.
  • $P_{DAC}$ is DAC precision ($P_{DAC}=1$ for bit-serial).
  • $N_{in}$ is the number of input cycles $\lceil b_{in}/P_{DAC}\rceil$
  • $b_{w}$ is the bit significance of the weights.
  • $b_{cell}$ is the number of bits per weight cell (e.g. $b_{w}=4$ for 4-bit MLC)
  • $N_{cell}$ is the number of cells per weight $\lceil b_{w}/b_{cell}\rceil$
  • $W_{i}$ represents the $i$-th bit slice of the weight matrix
  • $x_{j}$ represents the $j$-th bit slice of the input vector
For example, when using 4-bit MLC ($b_{cell}=4$) and 8-bit weights ($b_{w}=8$), $N_{cell}=2$ so $i$ ranges from 0 to 1 and the weight scales for each iteration are $2^{i\cdot b_{cell}}=$ $12^{0\cdot 4}=$ 1 and $2^{i\cdot b_{cell}}=$ $162^{1\cdot 4}=$ 16. When using bit-serial mode ($P_{DAC}=1$) to represent 8-bit activations ($b_{in}=8$), $N_{in}=8$ so $j$ ranges from 0 to 7 with input scales increasing by powers of 2. Each sum in equation 3 corresponds to a for-loop in the behavioral simulation. Consequently, the nested loop becomes the critical piece of code to optimize. Using PyTorch’s GPU-accelerated tensor operations, we can calculate the output $y$ of every memory array on the CIM chip design in parallel, which greatly reduces the overall runtime of NeuroSim.
Shown in Fig. 3, in device expert mode, three primary categories of device variations are modeled:
Device-to-Device (D2D) Variation: For each conductance (or capacitance) state $i$, the programmed value follows a Gaussian distribution:
$G=N(G_{mean_{i}},\sigma_{i})$ (4)
where $G_{mean_{i}}$ is the target conductance for state $i$ and $\sigma_{i}$ is the user-specified standard deviation. Separate variation parameters are specified for each level $i$. Users provide these parameters as comma-separated tuples, with one variation value per memory state. Users store these in the file ‘mem_states.csv’ that is then read by the framework automatically.
Stuck-at-Faults (SAF): Memory cells can become permanently fixed to their minimum or maximum states due to manufacturing defects or programming failures. This is modeled by randomly setting cells to minimum or maximum states according to user-specified probabilities before weight mapping. Unlike temporal variations, this permanent modification is applied once during initialization.
Temporal Drift: Time-dependent conductance changes follow:
$G(t)=G_{0}(t/t_{0})^{v}$ (5)
where $G_{0}$ is initial conductance, $t$ is retention time, $v$ is the drift coefficient, and $t_{0}$ is a reference time. The framework supports two drift modes: random drifting, or drifting towards a fixed state. The simulation treats arrays as frozen at the specified time point $t$ for inference calculations.
In circuit expert mode, statistical models are derived from Monte Carlo circuit analysis or silicon measurements. For SPICE-based characterization, users should:
  • Create a complete array model including peripheral circuits in a circuit simulator.
  • Run Monte Carlo simulations incorporating device variations (threshold voltage, device sizing), circuit-level effects (thermal noise, parasitic), and ADC quantization.
  • Generate statistical distributions by collecting multiple simulation results for each possible MAC output value.
For silicon-based characterization:
  1. 1.
    Collect repeated measurements across multiple array columns.
  2. 2.
    Compare expected vs. actual outputs over thousands of trials.
  3. 3.
    Build output distributions capturing combined effects of all noise sources.
The resulting models provide mean and standard deviation pairs for each possible MAC output value. The data can be displayed in a ”Confusion Matrix” like those shown in Fig. 3. Users provide these parameters as comma-separated tuples, with one variation value per output level. Users store these in the file ’output_noise.csv’. During behavioral simulation, the critical loop (Eq. 3 can be skipped, as the circuit behavior is modeled by the statistical output distribution. Therefore, we can calculate ideal partial sums with negligible runtime overhead. After ideal partial sums are calculated, the framework samples from the MAC distributions to efficiently inject noise into the outputs. This statistical approach enables rapid evaluation and SPICE-level accuracy through empirically validated noise characteristics (see case studies in Section IV.D).
TABLE I: Design Space Exploration Parameters
Parameter Range / Values
Technology Node 22nm
Device RRAM
[…]
Network / Dataset ResNet-18 / CIFAR-100
Array Dimensions $32\times 32$ $256\times 256$
Input / Weight Precision 6b, 8b
Input Processing bit-serial
Memory Cell Precision 1b -- 4b
ADC Precision 3b -- 11b
  • * The RRAM parameters are configured based on Intel’s 22nm FinFET platform [6]
III-D Performance Analysis Framework
To enable accurate power and performance estimation, NeuroSim V1.5 implements a trace-based interface between its behavioral simulation and hardware evaluation components. During inference simulation in PyTorch, the framework saves quantized input and weight values as decimal traces in CSV format (shown in Fig. 1. These traces capture the actual data patterns that would flow through the hardware, accounting for all quantization effects and bit slicing operations.
The hardware estimator (written in C++) automatically generates a floorplan for the entire accelerator based on the network architecture and user-specified parameters. It creates detailed transistor-level circuit models for all components (memory arrays, ADCs, interconnects, peripherals, accumulators, etc.), calibrated using SPICE simulations and data from previous tapeouts [7]. These models enable accurate hardware performance estimation at advanced technology nodes. Using traces from the software component, the estimator calculates detailed energy consumption for each array. The bit-serial input scheme is explicitly modeled - each input value is processed one bit at a time, with the framework tracking energy consumption for each operation. Weight values from the traces are converted to corresponding conductance or capacitance states based on parameters specified in the hardware configuration file.
In NeuroSim V1.5, we extended the hardware analyzer to support charge-based computation with non-volatile capacitive memories. This required the addition of new circuit models for nvCap arrays, charge-to-voltage conversion using an amplifier [61], and control circuits for the unique charge-based computation.
The hardware estimator uses these models along with device parameters specified in the configuration file to accurately evaluate PPA metrics for nvCap-based accelerators while maintaining compatibility with the existing trace-based interface between behavioral simulation and hardware analysis.
III-E Evaluation Setup
Our experimental evaluation consists of three main components: design space exploration of hardware configurations, inference accuracy analysis under various device non-idealities, and runtime performance comparison across open-source platforms. The DNN algorithms and proposed NeuroSim framework are implemented using PyTorch, with all simulations performed on a single NVIDIA RTX A6000 GPU with 48GB GDDR6.
III-E1 Design Space Exploration
For the design space exploration, we examine the trade-offs between hardware efficiency and inference accuracy across a range of architectural parameters. Table I summarizes the key parameters and their ranges used in the design space exploration study. The memory device characteristics are based on Intel’s 22nm RRAM platform [6]. We process inputs using the bit-serial approach (1b per cycle) to maintain consistency with the hardware estimator and activate all rows in parallel to maximize throughput.
The ADC precision requirements are determined by the dynamic range of array outputs, which depends on three key factors: memory cell precision ($b_{cell}$), DAC precision ($P_{DAC}$), and array rows ($R$). For a given array configuration, the maximum output value can be calculated as:
$out_{max}=R(2^{P_{DAC}}-1)(2^{b_{c}}-1)$ (6)
The minimum required ADC precision ($P_{ADC}$) to capture this full range without quantization loss is:
$P_{ADC}=\lceil log_{2}(out_{max})\rceil$ (7)
In our design space exploration, we use binary inputs (1b DAC) and we evaluate two ADC precision settings for each configuration:
  1. 1.
    Full precision matching the calculated dynamic range ($P_{ADC}$).
  2. 2.
    Reduced precision ($P_{ADC}-1$ and $P_{ADC}-2$) to study the impact of quantization noise.
Excluding device and circuit noise, the inference accuracy is solely affected by the input/weight quantization from floating-point to fixed-point representation and the quantization of array outputs depending on the ADC precision. Design parameters were swept using automated scripts to explore the trade-offs in accuracy and performance considering input/weight/ADC quantization and array size.
When reducing ADC precision below the ideal precision, we adopt a clipping approach, where the sensing margins for each analog output state remain the same regardless of ADC precision, and outputs larger than the precision of the ADC are clipped to the maximum ADC output based on its precision. We find that this method results in comparable accuracy to other, non-linear quantization methods while being the most practical to implement in hardware.
It is important to note that higher precision ADCs require smaller sensing margins between levels, which in practice could lead to increased susceptibility to circuit noise and higher error rates. The design-space analysis focuses solely on quantization effects, as detailed circuit-level noise characterization data is not available for the multi-bit memory configurations under study. For a complete understanding of system-level reliability, future work should incorporate both quantization effects and circuit-level noise measurements.
III-E2 Device Noise Case Studies
For analyzing inference accuracy under device non-idealities, we standardize the hardware configuration to isolate the effects of various noise sources and variations. Table II presents the fixed parameters used across all device non-ideality studies.
TABLE II: Device Non-idealities Case Study Parameters
Parameter Value
Technology Node 22nm
Device RRAM
[…]
Networks / Datasets VGG8 / CIFAR-10
ResNet-18 / CIFAR-100
ResNet-50 / ImageNet 1k
Swin-T / ImageNet 1k
Array Dimensions $128\times 128$
Activated rows 128
Input / Weight Precision 8b / 8b
Input Processing bit-serial
Memory Cell Precision 1b - 4b
ADC Precision* full precision
Device Noise Sources** D2D, SAF, state drift
  • * ADC precision is calculated using equations (6) and (7)
In these case studies, we analyze different sources of device variation including device-to-device variation, stuck-at-faults, and state drift across multiple DNNs and datasets to demonstrate support for multiple networks and noise types.
III-E3 Circuit Noise Case Studies
Next, we demonstrate the power of our improved noise modeling by using circuit noise data characterized both from SPICE characterization and tapeout measurements. With this improved modeling, direct comparisons between different devices, compute circuits, and DNNs can be performed. The table in Fig. 8 summarizes the key characteristics of each platform, with parameters drawn directly from their respective publications.
Mean and standard deviation for each output state are supplied in a comma separated format (.csv file), where each row of the file corresponds to the mean for each given output (post-ADC) and the standard deviation of the output.
III-E4 Runtime Evaluation
To characterize the improvements in runtime over previous versions of NeuroSim, we record the runtime of the inference accuracy portion under several key conditions. The main parameters that influence the runtime are the size of the network and dataset used the MLC level, precision of the DAC if one is used, and the input and weight precision.
IV Results and Case Studies
IV-A Design Space Exploration
We demonstrate NeuroSim V1.5’s capability for comprehensive design space exploration using a 22nm RRAM [6] platform as our baseline technology. Following the parameters outlined in Table I, we evaluate the fundamental tradeoffs between inference accuracy and hardware efficiency across architectural choices including array size, precision, and data representation. Our analysis focuses on architectural optimization without device/circuit noise effects, as these effects are technology-specific and addressed separately in subsequent noise modeling studies.
TABLE III: Default 22nm RRAM PPA (ResNet18 / CIFAR100)
TOPS TOPS/mm2 TOPS/W FPS Accuracy
11.6 0.013 21.3 7770 75.57%
  • Example performance using 22nm RRAM [6] and default settings in NeuroSim: $128\times 128$ array size, 7-bit ADC, 8b/8b input/weight precision.
Figure 4: Example design space exploration for 22nm RRAM. Each point represents a unique configuration. Bit-serial input was used for every design. Designs with highest TOPS/W use array sizes of $32\times 32$ and $64\times 64$.
We provide example performance metrics for the default configuration in NeuroSim using the selected RRAM device. As shown in Table III, we achieve 11.6 TOPS at 21.3 TOPS/W while processing 7,770 frames per second for CIFAR-100 images. Through systematic exploration of the design space, we can find designs that significantly outperform this baseline. Results are shown in Fig. 4. From this exploration, we identify several key trends that characterize optimal CIM architectures:
  1. 1.
    ADC Precision: The Pareto-optimal designs cluster in the 5--8 bit ADC precision range. Our analysis shows that ADC precision can generally be reduced by 1-bit from the lossless precision without significant accuracy degradation, and by 2-bits in certain configurations.
  2. 2.
    Array Dimensions: Optimal designs utilize array sizes between $32\times 32$ and $128\times 128$ cells. Smaller arrays reduce ADC precision requirements but necessitate additional accumulation circuitry, while larger arrays necessitate larger ADCs, incurring costs in latency and energy. The designs that achieve $>$40 TOPS/W used array sizes of either $32\times 32$ or $64\times 64$.
  3. 3.
    Memory Cell Precision: 2-bit and 3-bit MLC configurations dominate the Pareto frontier, offering an effective balance between storage density and programming reliability. These configurations achieve the highest energy efficiency (TOPS/W) for any given accuracy target.
The interplay between these parameters creates distinct architectural sweet spots. For example, the 8b/8b/64/6/2 configuration (input/weight/array size/ADC/MLC) achieves 75% accuracy while maintaining high energy efficiency. Designs with reduced precision (6b/6b) can push efficiency above 40 TOPS/W at the cost of $\sim 5\%$ accuracy degradation. The ability to rapidly evaluate these configurations through NeuroSim V1.5 enables the optimization of architectural parameters based on specific device and circuit designs.
Next, we demonstrate the improved noise-modeling capabilities in the new behavioral simulator.
IV-B Noise Modeling Case Studies
Figure 5: Inference accuracy of VGG8 on CIFAR-10, ResNet18 on CIFAR-100, and ResNet50 on ImageNet using V1.5 across different RRAM (single-level cell) variations.
IV-B1 D2D Variation
We evaluated inference accuracy across multiple neural networks using NeuroSim V1.5’s comprehensive noise modeling capabilities, including D2D variation, memory state drifting, stuck-at-faults and circuit-level MAC output noise. To systematically assess accuracy degradation, we selected network architectures of increasing complexity: VGG8 for CIFAR-10, ResNet18 for CIFAR-100, ResNet50 for ImageNet, and Swin Transformer Tiny for ImageNet. Following established methodology [62], we excluded the first convolutional layer from quantization in all simulations.
Figure 5 reveals that accuracy degradation becomes more severe with increasing network complexity and dataset difficulty. VGG8 on CIFAR-10 maintains reasonable accuracy until LRS variation reaches 20%. In contrast, ResNet18 on CIFAR-100 and ResNet50 on ImageNet show significant accuracy drops at just 10% and 5% LRS variation, respectively. Most notably, Swin Transformer Tiny demonstrates the lowest tolerance to device-level variations, highlighting the need for additional research into improving transformer model robustness for analog computing implementations.
Figure 6: Inference accuracy of VGG8 on CIFAR-10 regarding random device conductance drifting, drifting to Gmax, and drifting to Gmin with 1b cells.
IV-B2 Conductance Drift
Fig. 6 examines how three different types of conductance drift (modeled by Equation 5) affect inference accuracy. First, random drift shows better accuracy retention compared to drifting toward minimum or maximum conductance. This improved resilience occurs because cells at minimum conductance cannot drift lower, while cells at maximum conductance cannot drift higher, effectively limiting the total conductance change possible under random drift. Second, drifting toward maximum conductance maintains better accuracy than drifting toward minimum conductance. This asymmetry arises from Equation 5, where the drift coefficient v determines the drift direction. For the same retention time t and drifting speed, drifting toward minimum conductance results in a larger conductance difference compared to drifting toward maximum, leading to a larger deviation of MAC outputs.
Figure 7: Inference accuracy of VGG8 and ResNet18 under different stuck-at-fault rates.
IV-B3 Stuck-at-Faults
NeuroSim V1.5 adds modeling capabilities for stuck-at-fault (SAF) errors, extending beyond the D2D variation and drift modeling present in V1.4. Fig. 7 demonstrates how SAF errors impact inference accuracy across different networks, with SAF rates bounded by realistic RRAM fabrication constraints: 1.75% for LRS and 9.0% for HRS [42]. Our analysis reveals several key findings. First, both 1-bit and 2-bit cells show similar accuracy degradation at equivalent SAF rates, suggesting that cell precision does not significantly influence SAF sensitivity. However, network complexity strongly affects SAF tolerance - ResNet-50 on ImageNet loses accuracy even at minimal SAF rates, while VGG8 and ResNet18 maintain limited functionality only at the lowest SAF rates. Notably, SAF errors cause more severe accuracy degradation compared to other noise sources like D2D variation or conductance drift. This heightened sensitivity to SAF errors aligns with previous research findings [42], highlighting the critical importance of minimizing stuck faults during device fabrication.
Figure 8: (a) Actual MAC output with 2b FeFET current domain CIM [43] under 50mV Vth variation; (b) Actual MAC output with 2b FeFET charge domain CIM [43] under 50mV Vth variation; (c) Actual MAC output of RRAM CIM tape-out [41]; (d) Actual MAC output with 1b nvCap charge domain CIM [33, 18] under 50mV Vth variation; (e) Inference accuracy of various neural network algorithms across CIM macro A-D.
IV-B4 MAC Output Noise
circuit-level noise sources during array computation can be captured as statistical variations in the post-ADC MAC outputs. To demonstrate NeuroSim V1.5’s flexible noise modeling capabilities, we evaluated inference accuracy across four different CIM macros, incorporating both SPICE-characterized circuit data and silicon measurements from fabricated chips.
CIM A and B [43] employ 2b FeFET technology with different computing modes: CIM A performs current-mode computation with FeFETs directly, while CIM B implements charge-based computation using FeFET-modulated capacitors. Their noise characteristics were derived from Monte-Carlo SPICE simulations. CIM C [7] provides real silicon validation through direct measurements from a fabricated RRAM-based CIM chip. For these three designs, each MAC output level is characterized by unique mean and standard deviation values. CIM D [33, 18] uses 28 nm nvCap with charge-mode computation, where SPICE simulations revealed thermal noise as the dominant source, here, a uniform variation across output levels was used.
Figs. 8(a-d) illustrate the statistical distribution of actual versus ideal MAC outputs for representative cases. While CIM A, B, and D show tighter error bounds compared to CIM C, all designs achieve similar accuracy on smaller networks. Fig. 8(e) compares inference accuracy across architectures of increasing complexity. VGG8 and ResNet18 maintain reasonable accuracy across all implementations, with CIM C showing slightly higher degradation ( 2%) due to increased output variation. For ResNet50, only CIM B and D - which achieve lower standard deviations in their MAC errors - maintain high accuracy. The Swin Transformer exhibits particular sensitivity to output noise, with only the nvCap-based CIM D achieving software-comparable performance.
These results validate NeuroSim V1.5’s noise modeling framework as a versatile platform for evaluating CIM designs using either simulation data or silicon measurements. By providing a standardized methodology for capturing MAC output statistics, the framework enables direct comparison of different technologies and compute circuits under realistic operating conditions.
IV-C Runtime Evaluation and Simulator Comparison
We first compare NeuroSim V1.5 against existing open-source ACIM simulation frameworks to highlight key capabilities and performance improvements (Table IV). CrossSim [37] provides detailed device-level accuracy simulation but faces significant runtime overhead when modeling noise effects. AIHWKit [63] specializes in noise-aware training but offers limited support for general noise modeling and architectural exploration. While NeuroSim V1.4 [54] includes PPA analysis and has some noise modeling, it was constrained by slower inference speeds and limited network support.
TABLE IV: Comparison of DNN ACIM Simulators
DNN CIM Simulator
CrossSim
[37]
AIHWKit
[63]
NeuroSim
V1.4 [54]
NeuroSim
V1.5
Supported devices
PCM,
RRAM,
EcRAM,
DRAM
RRAM,
PCM,
Flash
SRAM,
RRAM,
FeFET
SRAM,
RRAM,
FeFET,
nvCap
Supported network types*
Linear,
Conv
Linear,
Conv,
Recurrent,
Transformer
Linear,
Conv,
Recurrent
Linear,
Conv,
Recurrent,
Transformer
Custom Compute Circuit Support No No No Yes
Bit Slicing Support Yes No Yes Yes
PPA Support No No Yes Yes
  • * PPA evaluation excluded.
The runtime is primarily influenced by the precision parameters that determine computation complexity through Equation 3: DAC precision affects the number of input cycles, cell bit-precision (MLC) determines the number of weight components, and noise modeling adds statistical sampling overhead. With 8b/8b/None configuration (DAC / MLC / Noise) shown in Table V, V1.5 achieves a baseline runtime of 0.95 ms/image for VGG8 and CIFAR-10. When using 1b cells to represent 8b weights, runtime increases $\sim$4$\times$ to 4.1 ms/image due to additional bit-slicing operations. Adding device noise results in negligible runtime increase and adding per-MAC output noise increases runtime to only 5.1 ms/image.
NeuroSim V1.5 achieves this dramatic improvement in simulation efficiency through three key advances: a PyTorch-optimized ACIM behavioral simulator, noise modeling with pre-characterized statistical noise modeling and TensorRT integration for efficient quantization. As shown in Fig. 9, these optimizations deliver a 6.51$\times$ speedup over V1.4 for CIFAR-10 inference with 1-bit cells.
Figure 9: Inference runtime of different neural network algorithms using V1.4 and V1.5 across 1b/2b/4b RRAM cell states without device non-idealities and noise.
Tables V and VI further demonstrate V1.5’s efficient noise modeling capabilities on larger networks towards CIFAR-100 and ImageNet datasets. Compared to CrossSim, which shows a significant slowdown when incorporating device noise, V1.5’s statistical approach adds minimal overhead - just 1.3$\times$ for uniform noise and 3.1$\times$ for output-dependent noise simulation. This efficiency stems from our pre-characterized noise distribution approach, eliminating the need for detailed circuit simulation during inference discussed in section III-B.
TABLE V: Inference Runtime (Seconds/Image) Comparison of VGG8 on CIFAR-10 (8b activation, 8b weight) between NeuroSim V1.4 and NeuroSim V1.5
DAC Precision/
/MLC/Noise
NeuroSim
V1.4 [54]
NeuroSim
V1.5
8b/8b/None 0.00095
8b/1b/None 0.0041
1b/8b/None 0.042 0.0048
1b/1b/None 0.17 0.027
8b/8b/Device Noise 0.00095
8b/1b/Device Noise 0.0041
1b/8b/Device Noise 0.042 0.0048
1b/1b/Device Noise 0.17 0.027
8b/8b/Output noise 0.0011
8b/1b/Output noise 0.0051
1b/8b/Output noise 0.0059
1b/1b/Output noise 0.039,0.085*
  • $128\times 128$ array size
  • * 0.039: same noise on each MAC output, 0.085: individual noise on each MAC output
These advances establish NeuroSim V1.5 as a comprehensive development platform that bridges detailed device/circuit-level analysis with system-level architectural exploration. The framework’s efficiency and adaptability make it particularly valuable as the field moves toward more complex network architectures requiring extensive accuracy and performance optimization.
TABLE VI: Inference Runtime (Seconds/Image) Comparison of ResNet50 on ImageNet (8b activation, 8b weight) between CrossSim, AIHWKit, and NeuroSim V1.5
DAC Precision/
/MLC/Noise
CrossSim
[37]
AIHWKit
[63]
NeuroSim
V1.5
8b/8b/None 0.3 0.0025 0.0085
8b/1b/None 1.1 0.033
1b/8b/None 1 0.04
1b/1b/None 5.8 0.24
8b/8b/Device Noise 9.8 0.003 0.0085
8b/1b/Device Noise 200 0.033
1b/8b/Device Noise 45 0.041
1b/1b/Device Noise 1220 0.24
8b/8b/Output noise 0.0025 0.0093
8b/1b/Output noise 0.039
1b/8b/Output noise 0.047
1b/1b/Output noise 0.29,0.61*
  • $128\times 128$ array size
  • * 0.29: same noise on each MAC output, 0.61: individual noise on each MAC output
V Conclusion
The growing complexity of AI models places increasing demands on CIM accelerator design, requiring tools that can efficiently evaluate both accuracy and hardware implementation trade-offs. This work presents NeuroSim V1.5, which combines accuracy analysis with power, performance, and area estimation to support systematic design space exploration of CIM architectures.
The framework’s integration with TensorRT enables evaluation of pre-trained networks without modification, while our flexible noise modeling methodology supports both detailed device-level analysis and efficient MAC-level noise injection, validated against published SPICE simulations and silicon measurements. These capabilities, combined with up to 6.5× faster runtime compared to NeuroSim V1.4, enable more extensive design space exploration than previous versions.
Through our case studies with a variety of device technologies and neural network topologies, we demonstrate how NeuroSim V1.5 can provide insights into CIM design trade-offs:
  1. 1.
    Optimal designs must balance array size, ADC precision, and multi-level cell capability.
  2. 2.
    Network complexity influences sensitivity to device-level variations, particularly in larger networks like ResNet-50 and Swin-T.
  3. 3.
    Charge-domain computation (e.g., based on nvCap) is a promising device for future large-scale ACIM acceleration.
Looking forward, the emergence of large language models (LLMs) with billions to trillions of parameters presents new challenges beyond on-chip memory capacity. These models will require GB-level high bandwidth memory (HBM) and process-in-memory (PIM) architectures that place compute logic near DRAM dies. While direct LLM and PIM support is beyond NeuroSim V1.5’s current scope, we have initiated efforts toward these capabilities through our work on 3D-stacked DRAM on TPU-like architectures [64, 65].

### v2

@@
 Abstract
-The exponential growth of artificial intelligence (AI) applications has exposed the inefficiency of conventional von Neumann architectures, where frequent data transfers between compute units and memory create significant energy and latency bottlenecks.
-Analog Computing-in-Memory (ACIM) addresses this challenge by performing multiply-accumulate (MAC) operations directly in the memory arrays, substantially reducing data movement.
-However, designing robust ACIM accelerators requires accurate modeling of device- and circuit-level non-idealities.
-In this work, we present NeuroSim V1.5, introducing several key advances: (1) seamless integration with TensorRT’s post-training quantization flow enabling support for more neural networks including transformers, (2) a flexible noise injection methodology built on pre-characterized statistical models, making it straightforward to incorporate data from SPICE simulations or silicon measurements, (3) expanded device support including emerging non-volatile capacitive memories, and (4) up to 6.5× faster runtime than NeuroSim V1.4 through optimized behavioral simulation.
-The combination of these capabilities uniquely enables systematic design space exploration across both accuracy and hardware efficiency metrics.
-Through multiple case studies, we demonstrate optimization of critical design parameters while maintaining network accuracy.
-By bridging high-fidelity noise modeling with efficient simulation, NeuroSim V1.5 advances the design and validation of next-generation ACIM accelerators.
-All NeuroSim versions are available open-source at https://github.com/neurosim/NeuroSim.
-Index Terms:
-Compute-in-memory, AI accelerator, hardware/software co-design, open-source, benchmark
+[…]
+publicationid: pubid: 0000–0000/00\$00.00 © 2026 IEEE
 I Introduction
-The exponential growth in AI applications has exposed a critical challenge in energy-efficient computing.
-The fundamental limitation lies in the mismatch between the memory-centric workloads such as in machine learning algorithms and the processing-centric architecture in contemporary hardware platforms.
-This mismatch causes massive data movement between memory and processing units to consume more than five times the energy of computation [1].
-Figure 1: Overview of NeuroSim V1.5: This update revamped the software backbone to support arbitrary neural network models and provide a more intuitive noise simulation workflow tailored to device and circuit experts.
-NeuroSim V1.5 enables users to evaluate inference accuracy across diverse architectures, including custom convolutional neural networks and transformers, while simulating various device types and noise sources.
-The platform uniquely supports power, performance, and area (PPA) estimations, facilitating comprehensive system-level design space exploration.
-Key new features include optimized noise modeling, compatibility with advanced non-volatile memory technologies, and improved runtime performance.
-The challenges stem from the data-intensive nature of deep neural networks (DNNs), where a relatively small number of operations is performed per fetched data; this causes system performance to be limited by memory bandwidth, as noted by the roofline model [2].
-Compute-in-memory (CIM) has emerged as a promising paradigm to tackle this bottleneck by co-locating data storage and multiply-accumulate (MAC) operations in the same physical arrays, thereby reducing repeated weight fetches.
-In digital computing-in-memory (DCIM), memory arrays (typically made of modified SRAM bit cells) are augmented with digital multipliers and adder trees for partial-sum accumulation, delivering robust performance with the overhead of adder trees [3].
-By contrast, ACIM leverages the intrinsic electrical behavior of memory cells to carry out multiplications directly in the current or charge domain.
-Specifically, weight values are encoded as conductance (or capacitance), and when an input voltage is applied to a row, the resultant column current (or charge) naturally represents the multiplication of input and weight.
-Summing all such column outputs yields the MAC result in the analog domain and then being converted back to the digital domain through analog-to-digital converters (ADCs) [4, 5].
-A variety of emerging non-volatile memory (NVM) devices have been utilized for these ACIM arrays, including resistive random-access memory (RRAM) [6, 7, 8, 9], phase-change memory (PCM) [10, 11, 12], ferroelectric field-effect transistors (FeFET) [13, 14, 15], and more recently non-volatile capacitors (nvCap) [16, 17, 18].
-ACIM implementations can also rely on modified SRAM bit cells, tapping into current-based summation [19, 20, 21, 22, 23, 24], or incorporate on-chip capacitors to accumulate charge [25, 26] or using them for time-domain computing [27].
-Several of these memory cells are depicted in a generic CIM macro with peripheral circuits as shown in Fig.
-1.
-Over the past few years, both DCIM and ACIM have been demonstrated in silicon with memory capacities up to a few megabytes [28]—sufficient for many on-chip inference scenarios such as computer vision tasks involving millions of parameters.
-DCIM generally offers the robustness of digital processing, making it attractive for high-performance computing, whereas ACIM excels in energy efficiency, an essential requirement for power-constrained edge devices.
-ACIM accelerators can theoretically surpass 100 TOPS/W [3, 9, 19, 26]—far exceeding the 1–10 TOPS/W range of contemporary von Neumann platforms—largely because they eliminate most weight-reloading costs by performing the MAC operations directly within the memory arrays.
-In this paper, one operation is defined as an 8b×8b MAC.
-The key challenge in ACIM is to manage inherently noisy analog computations to maintain the inference accuracy.
-These noise sources span multiple abstraction layers.
-At the device level, fabrication imperfections lead to device-to-device (D2D) variations and stuck-at-faults.
-Memory states can also drift over time, degrading compute accuracy [29, 30, 31, 32].
-At the circuit level, thermal noise and process variations in peripheral circuits degrade sensing margins, resulting in errors in the ADC processes [33, 18].
-At the system level, partial sums may be quantized at a low bit precision, compounding errors as layer outputs are passed from one stage to another [34, 35, 36].
-Together, these effects can reduce the inference accuracy of DNNs run on ACIM chips.
-Consequently, modeling and ultimately mitigating these non-idealities is critical for the design of robust ACIM architectures.
-These multi-layered noise effects create complex interactions between device characteristics, circuit design choices, and neural network requirements.
-Therefore optimizing ACIM accelerators requires systematic exploration across this design space to balance inference accuracy with hardware efficiency.
-Such an exploration requires simulation frameworks that can evaluate both accuracy and hardware metrics while maintaining practical runtimes even for modern large-scale networks.
-Direct SPICE simulation of analog computing arrays is prohibitively expensive for large neural networks, making it impractical to evaluate broad parameter spaces with modern datasets.
-While tools like CrossSim [37] and AIHWKit [38] offer detailed device modeling and noise-aware training respectively, they lack integrated hardware analysis capabilities and require extensive engineering effort to support new devices or circuits.
-Other frameworks like CIMLoop [39] focus on hardware metrics without considering accuracy under device and circuit non-idealities.
-These limitations make it difficult to systematically explore and optimize ACIM designs across both software accuracy and hardware efficiency metrics.
-NeuroSim [40] addresses these challenges through an integrated framework that combines behavioral simulation of CIM operation with automated hardware analysis.
-Building on this foundation, we present NeuroSim V1.5, which introduces several key advances to enable comprehensive design space exploration:
+The rapid growth in AI model size has exposed the importance of energy-efficient processors.
+Memory access dominates power consumption in AI processing, consuming more than five times the energy of computation [1].
+For large models, weights must be loaded periodically from DRAM, shifting the bottleneck from compute capability to memory bandwidth [2].
+I-A Compute-in-Memory
+Compute-in-memory (CIM) addresses this bottleneck by co-locating multiply-accumulate (MAC) operations with data storage.
+Digital CIM (DCIM) augments SRAM arrays with multipliers and adder trees [3].
+Analog CIM (ACIM) leverages memory cell physics for current-, charge-, or time-domain computation, where weights encode as conductance or capacitance, and column currents/charges are summed to yield MAC results [4, 5].
+Various non-volatile memory (NVM) devices enable ACIM arrays, including RRAM [6, 7, 8, 9], PCM [10, 11], FeFET [12, 13, 14], Flash [15], and nvCap [16, 17, 18].
+SRAM-based implementations also exist for current-based [19, 20, 21] or charge-based [22, 23] computation.
+Both approaches present distinct trade-offs.
+ACIM with NVM can achieve over 100 TOPS/W [3, 9] and high-density, enabling entire models to be stored on-chip.
+However, NVM devices suffer from limited write endurance, slow programming, and analog noise from device variations.
+DCIM offers robust digital computation, unlimited write endurance, and compatibility with advanced CMOS nodes but incurs higher area overhead and lower throughput.
+These complementary characteristics motivate heterogeneous architectures: ACIM excels for weight-stationary layers where weights are programmed once, while DCIM suits write-intensive operations like attention in transformers.
+As transformers increasingly dominate AI workloads—from vision transformers (ViTs) to large language models—extending CIM frameworks to support these architectures becomes critical.
+I-B CIM Design Challenges
+Optimizing CIM accelerators requires navigating a multi-dimensional design space spanning device characteristics (memory technology, precision, on/off ratio), circuit parameters (array dimensions, ADC precision), and system-level choices (quantization schemes, data mapping).
+The central challenge of ACIM is managing inherently noisy analog computations.
+Device-level fabrication imperfections cause device-to-device (D2D) variations and stuck-at-faults [24, 25].
+Additionally, memory states drift over time [26], and circuit-level thermal noise and process variations degrade ADC sensing margins [27, 18].
+At the system level, quantized partial sums compound errors through sequential layers [28, 29].
+These multi-layered effects interact in complex ways, necessitating simulation frameworks that can rapidly evaluate both accuracy and hardware metrics across the design space.
+I-C Existing Tool Limitations
+Direct SPICE simulation is prohibitively slow for large networks, making broad parameter exploration impractical.
+Existing tools present different trade-offs: CrossSim [30] provides high-fidelity physical simulations but with high computational cost; AIHWKit [31] offers noise modeling primarily for training; CIMLoop [32] focuses on system architectures without accuracy modeling.
+No existing framework provides integrated accuracy-PPA co-optimization, particularly for trasnformers where self-attention mechanisms introduce unique challenges for CIM deployment.
+In this paper, we present NeuroSim V1.5, an integrated simulation framework for systematic co-optimization of accuracy and hardware efficiency in CIM accelerator design.
+Unlike tools that address device physics or system performance in isolation, NeuroSim V1.5 combines behavioral simulation with detailed hardware analysis to enable comprehensive design space exploration.
+Our key contributions are:
   1. 1.
-    Seamless integration with TensorRT to support arbitrary pre-trained models, while automating quantization.
+    Support for transformers with hybrid ACIM/DCIM architectures: Our hybrid design augments NeuroSim’s floorplan generator with DCIM tiles used for attention and LUTs for complex activation functions.
+We conduct a case study on ViT performance using this architecture and compare with CNNs.
   2. 2.
-    A flexible noise modeling methodology built on pre-characterized statistical models, enabling rapid evaluation of diverse device and circuit designs.
+    An integrated co-optimization framework combining TensorRT-based quantization, flexible noise modeling validated against SPICE/silicon data, and circuit-level PPA estimation—allowing rapid evaluation of accuracy-efficiency trade-offs across the full design space.
   3. 3.
-    Support for emerging technologies like charge-based non-volatile capacitive memories (nvCap) and emerging algorithms like vision transformer (ViT).
+    Expanded device support, including emerging non-volatile capacitive memories (nvCap) and charge-domain compute methods.
   4. 4.
-    Significantly improved runtime performance - up to 6.5× faster than previous versions of NeuroSim.
-These capabilities maintain seamless compatibility with NeuroSim’s existing power, performance, and area (PPA) analysis framework, creating a unique environment for systematic CIM accelerator design optimization.
-The framework achieves runtime only two to three times that of software-only standard GPU inference.
-The remainder of this paper is organized as follows: Section II provides technical background on ACIM systems and existing simulation approaches.
-Section III details our improved framework architecture and generic noise modeling methodology.
-Section IV presents results from various case studies including inference accuracy analysis of several devices and compute circuits using data from SPICE simulations and tapeout measurements [41, 42, 43, 33, 18, 7].
-We also provide design space exploration for a RRAM based ACIM chip to demonstrate the framework’s adaptability and improved running speed.
-Finally, Section VI discusses implications for future larger-scale AI model benchmarking and concludes the paper.
+[…]
+The remainder of this paper is organized as follows: Section II provides technical background on CIM systems.
+Section III details our framework architecture and noise modeling methodology.
+Section IV presents case studies including design space exploration, noise sensitivity analysis validated against SPICE and silicon measurements [33, 34, 27, 7], PPA breakdown of Swin Transformer on ImageNet, and detailed error analysis comparing CNN and ViT architectures.
+Section V discusses future directions.
 II Background Information
 II-A CIM Array Architecture and Operation
-ACIM performs matrix computations by utilizing physical properties of memory arrays.
-The fundamental building block is the crossbar array, where memory elements are arranged at the intersections of wordlines (WLs) and bitlines (BLs).
-This crossbar arrangement is uniquely suited for parallel MAC operations because each row’s input signal (voltage) is simultaneously applied to all memory devices along that row.
-The resulting outputs (currents or charges) are summed along each column, naturally implementing the dot product.
-Consequently, large-scale MAC operations can be performed in one analog ‘step’, vastly increasing the throughput compared to digital architectures that must shuttle data back and forth between separate memory and logic units.
-To date, two primary methods of analog computation have emerged.
-Resistive arrays constructed with RRAM, PCM, FeFET that encode neural network weights as conductance states (G), performing MAC operations in the current domain:
+ACIM performs matrix computations by utilizing electrical properties of memory arrays.
+The fundamental building block is the crossbar array, where memory elements are placed at wordline (WL) and bitline (BL) intersections.
+Crossbar arrays enable parallel MACs: input voltages applied to rows produce column currents/charges that naturally sum to implement dot products.
+This allows large-scale MAC operations in one analog step, vastly increasing throughput.
+Two primary analog computation methods have emerged.
+Resistor-based arrays contain RRAM, PCM, Flash, FeFET or other resistive memory, which encode neural network weights as conductance states (G) and perform MAC operations in the current-domain:
 $I_{BL_{i}}=\sum_{j}G_{ij}V_{j}$ (1)
-where $I_{BL_{i}}$ represents the analog matrix-vector multiplication (MVM) output current on the $i^{th}$ bitline, and $j$ indexes across the wordlines.
-Capacitive arrays utilize non-volatile capacitors (nvCap), where the weight is represented by the programmable capacitance of the ferroelectric or charge-trap memory cell [44] [45], thus performing MAC operations in the charge domain:
+where $I_{BL_{i}}$ is the analog matrix-vector multiplication (MVM) output current on the $i^{th}$ bitline, and $j$ indexes across wordlines.
+Capacitor-based arrays contain non-volatile capacitors (nvCap), which encode weights as programmable capacitance [35, 36], performing MAC operations in the charge domain:
 $Q_{BL_{i}}=\sum_{j}C_{ij}V_{j}$ (2)
-Alternative charge-based computation approaches utilize fixed capacitors in conjunction with SRAM cells, where the stored digital weight modulates charge accumulation through charge sharing mechanisms [46, 47].
-DCIM has emerged as another promising approach [3], where SRAM arrays are augmented with digital multipliers and adder trees to perform integer MAC operations directly within memory.
-These systems are attractive due to their compatibility with advanced CMOS technology scaling and robust digital operation.
-However, the digital processing overhead - particularly from the complex adder trees required for partial sum accumulation - impacts array density.
-Additionally, SRAM leakage power becomes more significant in advanced technology nodes.
-Both the array scalability and leakage power create challenges for running complete neural networks on-chip.
-Each approach is finding its niche, while ACIM remains near the theoretical limit of efficiency for neural network applications.
-In ACIM designs, the choice of memory technology influences key array characteristics like precision, reliability, and energy efficiency - topics we will explore in detail when discussing noise modeling methodology.
+Alternative charge-based computation approaches use fixed capacitors with SRAM cells, where stored digital weight modulates charge accumulation via charge sharing [37, 38].
+DCIM [3] uses digital multipliers and adder trees for in-memory integer MACs, offering robustness at the cost of higher area overhead compared to ACIM at the same technology node.
+The main advantage of DCIM is that it can be integrated in any CMOS technology, whereas non-volatile memories are mostly limited to legacy nodes.
 II-B Device and Circuit Level Considerations
-The precision and reliability of analog computing operations depend heavily on the properties of individual memory cells. Key device parameters include:
+The precision and reliability of analog computing operations depend heavily on individual memory cell properties. Key device parameters include:
 On/Off Ratio: The ratio between maximum and minimum conductance/capacitance states defines the usable dynamic range for weight storage.
-Finite on/off ratios introduce non-zero current/charge contribution from cells programmed to the off states, which accumulates across array columns and can degrade computation accuracy.
-This effect becomes particularly significant in larger arrays where hundreds of devices in the off-state contribute to the output.
-To address this, array designs typically include additional ‘dummy’ columns programmed to off states and are supported in NeuroSim [40], whose outputs are subtracted from the computational columns to cancel the accumulated off-state current/charge.
-While this technique effectively eliminates the impact of finite on/off ratio, it slightly reduces the available dynamic range for computation since part of the range must be reserved for cancellation.
-This results in smaller sensing margins and potentially smaller signal-to-noise ratio (SNR).
-Therefore, high on/off ratios are still a desired characteristic for ACIM memory devices.
+Finite on/off ratios introduce non-zero current/charge from off-state cells that accumulates across array columns and can degrade accuracy, especially in larger arrays.
+To address this, designs often include ‘dummy’ columns (supported in NeuroSim [39]) whose outputs are subtracted to cancel accumulated off-state current/charge.
+While effective, this slightly reduces computation dynamic range, resulting in smaller sensing margins and potentially smaller signal-to-noise ratio (SNR).
+Thus, high on/off ratios remain desirable for ACIM devices.
 Number of States: Supporting high-precision weights requires either multi-level memory cells (MLC) or bit slicing with multiple lower-precision/binary cells.
-For MLC devices, the number of reliably programmable states determines the weight precision.
-RRAM, PCM, and nvCap typically achieve 1-4 bits per cell [8, 28, 11], while FeFET have demonstrated >5 bits [14, 48].
-At the circuit level, array size is limited mainly by sensing margins, where reliable output detection requires sufficient separation between computational states.
-This becomes challenging with larger arrays due to increased noise accumulation and reduced signal levels from parasitic effects.
-Critical peripheral circuits include:
-  • WL/BL drivers to supply the input and output signals.
-  • ADCs for digitizing array outputs.
-  • Shift-and-add circuits for managing bit-precise computations.
-  • Accumulation units on the PE level for partial sum aggregation.
-  • Multiplexers for data routing and array access.
-The digital circuit blocks in the periphery can be verified through conventional VLSI design methodologies. Therefore, behavioral modeling efforts focus primarily on the analog computational elements.
-Figure 2: Procedure for ACIM behavioral simulation in NeuroSim V1.5.
+For MLC devices, the number of reliable programmable states determines weight precision.
+RRAM, PCM, Flash, and nvCap typically achieve 1-4 bits per cell [8, 40, 11], while FeFETs have demonstrated >>5 bits [13, 41].
+At the circuit level, array size is limited mainly by sensing margins, as reliable output detection requires sufficient separation between analog states.
+This becomes challenging with larger arrays due to increased noise accumulation and reduced signal levels from parasitics.
+Peripheral circuits include WL/BL drivers, ADCs, shift-and-add circuits, inter-array accumulation units, and multiplexers for data routing.
+Peripheral digital circuit blocks are verifiable via conventional VLSI methods; thus our behavioral modeling focuses primarily on analog computational elements.
+II-C Precision and Data Representation
+Neural network parameters must be carefully mapped to CIM arrays, managing precision and computational accuracy.
+This presents two key challenges: representing multi-bit input activations and storing high-precision weights.
+For input activations, three main approaches exist:
+  1. 1.
+    Multi-bit DACs convert inputs to analog voltages but require significant area/energy and may introduce device non-linearity errors.
+  2. 2.
+    Bit-serial processing eliminates DAC overhead at the cost of multiple computational cycles.
+  3. 3.
+    Time-encoding [42, 43, 44] offers simplified ADCs at the cost of latency. Time-domain operation is not currently supported in NeuroSim.
+For weight storage, two methods are available:
+  1. 1.
+    Bit Slicing: Used when desired weight precision ($b_{w}$) exceeds memory cell bit precision ($b_{cell}$). E.g., an 8-bit weight can be split across multiple memory cells.
+  2. 2.
+    Direct Mapping: Used when weight precision ($b_{w}$) is less than or equal to cell precision ($b_{cell}$).
+In this case, $N_{cell}=1$, and the entire analog weight maps directly to a single memory cell’s state.
+Bit-sliced weights occupy adjacent columns; outputs are scaled and summed post-ADC (’shift-add’, Fig.
+2).
+The PPA estimator models bit-serial inputs for hardware accuracy.
+Pre-ADC analog shift-add for multi-bit weights [45, 46] and pre-ADC input accumulation schemes [47, 48] can reduce compute cycles due to bit-slicing.
+While we have explored these techniques [35], they are not explicitly supported in this version but can be added with minimal user modification.
+CIM MACs operate in integer domain, requiring float-to-int quantization.
+ADCs digitize analog results, which are then de-quantized to floating-point.
+Some accelerators explore quantizing the entire network to integers [49, 50], requiring specialized approximations for operations like normalization and activation functions.
+[…]
+II-D Hierarchical System Architecture
+CIM accelerator organization stems from device technology and neural network structure constraints.
+Memory array sizes are typically $32\times 32$ to $256\times 256$ cells due to sensing margins and manufacturing yields.
+However, neural network layers often require much larger matrix operations, necessitating hierarchical decomposition.
+NeuroSim’s chip hierarchy: At base level, crossbar arrays compute partial sums of larger linear/convolutional layers.
+Crossbars are grouped into processing elements (PEs), where partial sums are accumulated.
+PEs are then grouped into tiles which include activation buffers and accumulation circuits for combining PE results– enabling efficient large layer processing by distributing computations while maintaining data locality.
+At chip level, multiple tiles operate via an H-tree or X-Y bus interconnect.
+This highest level includes global buffers for inter-tile data movement and accumulation circuits for distributed computation results.
+The system implements layer-wise pipelining, where different tiles process consecutive network layers simultaneously to maximize throughput.
+NeuroSim assumes all weights fit on-chip in NVM, enabling a fully weight-stationary dataflow.
+This design avoids frequent reprogramming that would stress NVM write endurance and increase latency.
+Small models that would underutilize a design’s compute capacity are duplicated to boost throughput [51].
+For models exceeding single-die capacity, we have developed separate frameworks for 3D chiplet integration [52].
+The PPA estimator generates hardware configurations optimized to a specific network architecture (e.g.
+CNNs).
+While the floorplan is generated based on one selected network architecture, the hardware can execute multiple models of the same type through runtime reconfigurable weight mapping and dataflow control, as demonstrated in our prior work [53].
+To simulate latency, area, and energy consumption, NeuroSim uses analytical models (C++) of circuit blocks based on standard cell libraries in advanced logic nodes.
+V1.4 [54] expanded technology support to 1nm, including stacked nanosheet transistors, as well as support for future A5 node with complementary-FETs.
+NeuroSim’s software/hardware co-design philosophy and open-source model enable research on emerging devices, array architectures, on-chip training [55], and advanced packaging [52, 56, 57].
+For a more detailed hierarchy analysis, see [39].
+Figure 2: Procedure for CIM behavioral simulation in NeuroSim V1.5.
 (1) Inputs and weights are quantized from floating point to integer.
 (2) Inputs and weights unfolded and (3) mapped to compute-in-memory arrays.
-(4) Analog MACs are calculated with analytical circuit models.
-(5) Noise is injected either to memory devices or to digital MAC outputs.
-(6) ADC outputs are shifted and added.
+(4) MACs are calculated with analytical circuit models.
+(5) Noise is injected either to memory devices (pre-ADC) or to digital MAC outputs (post-ADC).
+(6) ADC/Adder tree outputs are shifted and added.
 (7) Outputs from memory arrays are accumulated and concatenated.
@@
 (9) Outputs are de-quantized from integer back to floating-point.
-II-C Precision and Data Representation
-Neural network parameters must be carefully mapped to CIM arrays while managing precision and computational accuracy.
-This mapping presents two key challenges: representing multi-bit input activations and storing high-precision weights.
-For input activations, two main approaches exist:
-  1. 1.
-    Using digital-to-analog converters (DACs) to represent each input value as an analog voltage.
-While straightforward, this requires complex multi-bit DACs for every row of the array, introducing significant area and energy overhead.
-The I-V or C-V nonlinearity of the underlying memory devices may also introduces the errors in the input representation.
-  2. 2.
-    Processing each bit of the digital input serially, with outputs shifted and added according to bit significance. This eliminates DAC overhead but requires multiple computational cycles.
-For weight storage, two methods are available:
-  1. 1.
-    Bit slicing across multiple lower-precision cells.
-For example, an 8-bit weight can be distributed across cells of various precisions - from a single 8-bit cell to eight 1-bit cells, with intermediate combinations possible.
-When slicing the weight among multiple cells, each cell is placed in an adjacent column, with the outputs of each column scaled by its significance post-ADC and added with the other scaled outputs.
-This operation is referred to as ‘shift-add’ and is described in Fig.
-2 and in section III.
-Using analog shift-add before the ADCs has also been explored [44, 49, 50], but is not explicitly supported in NeuroSim V1.5.
-While NeuroSim’s behavioral simulator supports both input representation methods, the PPA estimator only supports the bit-serial approach due to its practical hardware implementation.
-In CIM, the fundamental MAC operation is performed in the integer domain, necessitating quantization of the inputs and weights of each layer in the network from floating-point to integer.
-The analog multiplication and accumulation processes generate integer results, which are digitized through ADCs.
-Finally, the integer outputs are de-quantized back to floating-point for further processing.
-Some accelerators explore quantizing the entire network to integers[51, 52], requiring specialized approximations for operations like normalization and activation functions.
-This approach is particularly applicable to transformers, which require more complex operators like GELU and layer normalization.
-II-D Hierarchical System Architecture
-The organization of ACIM accelerators stems from fundamental constraints in both device technology and neural network structure.
-Memory array sizes are typically limited to dimensions between 32×32 and 256×256 cells due to sensing margins and manufacturing considerations.
-However, neural network layers often require much larger matrix operations.
-This dimensional mismatch necessitates a hierarchical decomposition of network operations.
-The organization of the chip hierarchy in NeuroSim is as follows: At the base level, synaptic arrays perform MAC operations through crossbar structures.
-These arrays are grouped into processing elements (PEs), which implement weight-stationary, spatial mapping of convolutional operations.
-For example, when mapping 3D convolutional kernels to 2D array structures, multiple arrays within a PE process different kernel depths in parallel, with their outputs appropriately accumulated.
-PEs are further aggregated into computational tiles, each capable of processing a portion of a neural network layer.
-The tile architecture incorporates buffers for activation storage and accumulation circuits for combining partial results from multiple PEs.
-This organization enables efficient processing of large layers by distributing computations across multiple tiles while maintaining data locality.
-At the chip level, multiple tiles operate together through either an H-tree or X-Y bus interconnect structure.
-This highest level of hierarchy includes global buffers for inter-tile data movement and accumulation circuits for combining results from distributed computations.
-The entire system implements layer-wise pipelining, where different tiles process consecutive layers of the network simultaneously to maximize throughput.
-To balance processing speed between layers of different sizes, weight duplication can be employed within tiles [53].
-For instance, if a layer’s weight matrix is smaller than the tile’s capacity, the weights can be duplicated to process multiple input vectors simultaneously, improving throughput.
-This hierarchical structure enables flexible mapping of diverse neural network architectures while managing the physical constraints of ACIM arrays.
-To model this hierarchy, NeuroSim employs analytical models for circuit blocks derived from standard cell libraries in advanced logic nodes.
-The previously released V1.4 [54] expanded technology support to the leading-edge 1 nm node, introducing stacked nanosheet transistors and studying the scaling path toward a future A5 node with complementary-FET (CFET) technology.
-Thanks to NeuroSim’s software/hardware co-design philosophy and open-source development model, the tool enables broader research on emerging device technologies, array architectures, on-chip training [55], and advanced packaging [56, 57, 58].
-For a detailed analysis of the hierarchy in NeuroSim, readers are referred to [40].
 III Simulation Methodology
 III-A Overview of Design Methodology
-Designing ACIM accelerators requires optimization across multiple abstraction layers - from device characteristics to system architectures.
-NeuroSim V1.5 provides an integrated framework to evaluate these designs through two main components shown in Fig.
-1 working jointly:
+Designing CIM accelerators requires optimization across multiple abstraction layers—from device characteristics to system architectures.
+NeuroSim V1.5 provides an integrated framework to evaluate these designs through two main components (Fig.
+1) working jointly:
   1. 1.
-    A behavioral simulator that maps and quantizes neural networks for ACIM implementation and evaluates inference accuracy while accounting for hardware non-idealities.
+    A behavioral simulator that maps quantized neural networks to the CIM floorplan and evaluates inference accuracy under hardware non-idealities.
   2. 2.
-    A hardware analyzer that estimates system-level energy, latency, and area through detailed circuit models of array peripherals and interconnects.
-In this update, our first contribution is a comprehensive re-design of the behavioral simulator.
-Second, we update the hardware analyzer to support the emerging nvCap-based CIM, complete with circuit models for the new charge-domain compute mode.
-In the behavioral simulator, we have fundamentally enhanced how we model compute-in-memory operations for accuracy estimation.
-A key approach is that we leverage pre-characterized statistical models derived from SPICE simulations or silicon measurements to ensure accuracy, while utilizing GPU-accelerated PyTorch operations to achieve the throughput needed for large-scale inference simulation.
-This separation allows us to maintain high fidelity while dramatically improving simulation efficiency.
-The behavioral simulator supports two primary simulation modes:
-Device expert mode: Model memory array physics directly, simulating conductance variations, stuck faults, and temporal drift to isolate the effect of memory technologies on network accuracy.
-Circuit expert mode: Model memory arrays with a statistical approach using pre-characterized noise distributions from SPICE simulations or silicon measurements, enabling rapid evaluation of aggregated non-idealities from devices and circuits.
-Through integration with TensorRT’s post-training quantization flow, NeuroSim V1.5 automatically maps pre-trained neural networks to the corresponding ACIM architecture.
-The following sections detail the quantization methodology, array modeling approach, and specific techniques for incorporating device and circuit-level non-idealities into the simulation framework.
+    A hardware analyzer that estimates system-level energy, latency, and area through detailed circuit models of arrays, peripherals, and interconnects.
+The behavioral simulator supports two primary simulation modes: Device expert mode: Model memory array physics directly (conductance variations, stuck faults, temporal drift) to isolate memory technology effects on network accuracy.
+Circuit expert mode: Model memory arrays statistically using pre-characterized noise distributions (SPICE/silicon), enabling rapid evaluation of aggregated device/circuit non-idealities.
+Through TensorRT’s post-training quantization, NeuroSim V1.5 automatically maps pre-trained neural networks to the corresponding CIM architecture.
+Subsequent sections detail quantization, array modeling, and incorporating non-idealities.
 III-B Neural Network Quantization and Mapping
 III-B1 Quantization
-NeuroSim V1.5 adopts NVIDIA’s TensorRT quantization infrastructure for converting neural networks from floating-point to integer representations.
-This differs from previous NeuroSim versions which used in-house WAGE quantization [59] - an integer-training method that requires retraining the network.
-TensorRT instead provides post-training quantization through automatic operator replacement, supporting quantization of pre-trained networks using operators like linear layers, convolutions, and pooling without modification.
-TensorRT supports configurable bit widths for both inputs and weights, which we use to evaluate different precision settings (e.g., 8b/8b, 6b/6b).
-The framework determines optimal scaling factors through either max or histogram calibration during a brief calibration phase.
-This phase typically requires only 2 batches of calibration data (e.g., several hundred images) to analyze the distribution of values in both input tensors and weight matrices.
-Max calibration simply uses the maximum observed value across these tensors to set the scaling factor.
-Histogram calibration creates a histogram of the observed tensor values and sets the scale factor based on the cumulative distribution function (CDF) - specifically, choosing the value at which the CDF reaches a user-specified percentile (typically 99.99%).
-This helps exclude outliers that would otherwise force a larger dynamic range.
-The key advantages of adopting TensorRT over WAGE quantization are direct support for arbitrary pre-trained models without retraining and the automated support for multiple operator types through operator replacement.
+NeuroSim V1.5 uses TensorRT’s post-training quantization, unlike previous WAGE training-based methods [58].
+TensorRT automatically replaces operators (linear, convolution, pooling) with configurable precision (e.g., 8b/8b, 6b/6b).
+Scaling factors use max or histogram calibration (99.99% CDF percentile, 2 batches).
 III-B2 Network to Array Mapping
-Physical constraints of ACIM arrays necessitate careful partitioning of neural network operations.
-For a linear layer with input dimension $N$ and output dimension $M$, the quantized weight matrix must be mapped to multiple arrays based on both the memory array dimensions and the precision requirements.
-The number of memory cells required per weight depends on the bit precision of the cells ($b_{cell}$).
-For example, to represent an 8-bit weight using 1-bit cells, we need $N_{cell}=\lceil 8/1\rceil=8$ cells per weight.
-This bit-slicing requirement effectively multiplies the width dimension of our mapping by $N_{cell}$.
-As shown for a linear layer in Fig.
-3, these weight components are arranged in adjacent columns, with each column corresponding to a different bit significance.
-Memory array size further constrains the mapping.
-With a memory array height of $R$ rows and $C$ columns, the $N\times M$ weight matrix must be partitioned into $\lceil N/R\rceil$ rows of arrays.
-The total number of arrays needed to map the weights is equal to $\lceil N/R\rceil\cdot\lceil MN_{cell}/C\rceil$.
-Mapping convolutional layers to memory arrays requires additional unfolding operations, of which multiple methods can be implemented.
-For a more detailed description, readers are referred to [40].
-The input feature map is unfolded and mapped to the arrays according to the selected input scheme.
-In Fig.
-2 we show the bit-serial operation, where the feature map is processed separately for each bit of the quantized input.
-Outputs from vertically-aligned arrays are accumulated, as these represent partial sums from the same output neurons.
-Outputs from horizontally aligned arrays are shifted and added according to their bit significance, then concatenated to form the complete layer output.
-III-B3 Support for Transformers
-For transformer architectures, NeuroSim V1.5 adopts a hybrid approach that maps compatible operations to CIM arrays while implementing other operations in the digital domain.
-The framework maps linear layers (including query, key, value projections and MLP layers) to CIM arrays using the same methodology as standard neural networks.
-These linear layers comprise a significant portion of the total computation in transformer models.
-However, dynamic operations such as attention score computation, layer normalization, and activation functions must still be computed.
-We elect to compute these dynamic operations in floating-point compute units due to CIM hardware constraints.
-Non-volatile memory-based CIM is optimal for weight-stationary computations that do not require frequent weight updates.
-This is due to limited write endurance and slow programming speed common in these devices.
-Therefore, operators like self-attention that require unique input and weight matrices each inference are not well-suited for non-volatile CIM.
-Digital CIM implementations using SRAM could potentially handle these dynamic operations, however they would still require careful quantization of the query, key, and value matrices to maintain compatibility with integer-only computation [60].
-Recent work on integer-only transformers [51, 52] has demonstrated promising approaches for full integer quantization, though further research is needed to evaluate the accuracy-efficiency tradeoffs of the hardware.
-The current hardware estimation portion of NeuroSim does not include circuit models for floating-point functional units needed for self-attention computation, layer normalization, or GELU activation functions.
-Future versions will explore integrating these components, potentially through integer approximations that maintain accuracy while enabling efficient hardware implementation.
-Figure 3: NeuroSim V1.5 flowchart.
+Physical CIM array constraints require careful neural network operation partitioning.
+For example, for a linear layer (input $N$, output $M$), the quantized weight matrix maps to multiple arrays based on array dimensions and precision.
+The number of memory cells per weight depends on cell bit precision ($b_{cell}$).
+E.g., an 8-bit weight using 1-bit cells needs $N_{cell}=\lceil 8/1\rceil=8$ cells.
+Shown for a linear layer in Fig.
+2, these weight components split among adjacent columns, each column representing a different bit significance.
+Memory array size (R rows, C columns) further constrains mapping.
+An $N\times M$ weight matrix partitions into $\lceil N/R\rceil$ array rows.
+Total arrays for weights: $\lceil N/R\rceil\cdot\lceil MN_{cell}/C\rceil$.
+To map 3D convolutional layers to 2D matrices we used im2col-like mapping described in [39].
+The input feature map unfolds and maps per selected input scheme.
+Fig.
+2 shows bit-serial operation (feature map processed separately per quantized input bit).
+Outputs from vertically-aligned arrays accumulate (partial sums for same output neurons).
+Outputs from horizontally-aligned arrays shift and add by bit significance, then concatenate for complete layer output.
+Figure 3: NeuroSim V1.5 noise modeling flowchart.
 Device expert mode: user provides device data (memory state variation, stuck-at-faults, state drift).
@@
 III-C1 Memory Array Operation
-Building on TensorRT’s operator replacement strategy, NeuroSim V1.5 substitutes standard neural network layers with specialized CIM operators that model array-level computation.
-When a PyTorch model is loaded, linear and convolutional layers are automatically replaced with cim.Linear and cim.Conv modules that partition operations according to array constraints and simulate analog computation as described in Fig.
-3.
-The framework incorporates two separate simulation modes for modeling non-idealities: device expert mode for measured device-level device variation or calibrated aging models, and circuit expert mode using statistical models derived from SPICE or silicon measurements.
-These modes cannot be used simultaneously as circuit-level noise inherently captures device-level variations.
+NeuroSim V1.5 replaces neural network layers with specialized CIM operators.
+Runtime improvements stem from: (1) optimized GPU-accelerated MVM simulation in ‘cim.Linear‘ and ‘cim.Conv‘ modules, and (2) efficient TensorRT quantization.
+Linear and convolutional layers auto-replace with custom modules that partition operations and simulate analog computation (Fig.
+3).
+Two simulation modes exist: device expert (measured variations, aging models) and circuit expert (SPICE/silicon statistics), used separately to avoid double-counting.
 III-C2 Device and Circuit Non-Idealities
-In device expert mode, integer weights are first mapped to conductance or capacitance states based on the target memory technology.
-The framework then simulates analog multiplication through Ohm’s law (I = GV) for resistive arrays or charge-based computation (Q = CV) for capacitive arrays.
-Input values are treated as idealized voltages, and outputs are computed by summing the resulting currents or charges along each column.
+In device expert mode, integer weights map to conductance/capacitance states.
+The framework simulates analog multiplication via Ohm’s law ($I=GV$) for resistive arrays or charge-based computation ($Q=CV$) for capacitive arrays.
+Inputs are idealized voltages; outputs are summed column currents/charges.
+Noise reflecting device non-idealities is typically injected at this stage, pre-ADC quantization.
 To capture all possible input and weight bit-slicing combinations, including both bit-serial and DAC-based input schemes, we use the following generalized equation:
@@
   • $x_{j}$ represents the $j$-th bit slice of the input vector
-For example, when using 4-bit MLC ($b_{cell}=4$) and 8-bit weights ($b_{w}=8$), $N_{cell}=2$ so $i$ ranges from 0 to 1 and the weight scales for each iteration are $2^{i\cdot b_{cell}}=$ $12^{0\cdot 4}=$ 1 and $2^{i\cdot b_{cell}}=$ $162^{1\cdot 4}=$ 16.
+For example, when using 4-bit MLC ($b_{cell}=4$) and 8-bit weights ($b_{w}=8$), $N_{cell}=2$ so $i$ ranges from 0 to 1 and the weight scales for each iteration are $2^{i\cdot b_{cell}}=$ $2^{0\cdot 4}=1$ and $2^{i\cdot b_{cell}}=$ $2^{1\cdot 4}=16$.
 When using bit-serial mode ($P_{DAC}=1$) to represent 8-bit activations ($b_{in}=8$), $N_{in}=8$ so $j$ ranges from 0 to 7 with input scales increasing by powers of 2.
@@
 Users store these in the file ‘mem_states.csv’ that is then read by the framework automatically.
-Stuck-at-Faults (SAF): Memory cells can become permanently fixed to their minimum or maximum states due to manufacturing defects or programming failures.
-This is modeled by randomly setting cells to minimum or maximum states according to user-specified probabilities before weight mapping.
-Unlike temporal variations, this permanent modification is applied once during initialization.
+Stuck-at-Faults (SAF): Cells permanently fixed at min/max states due to defects. Modeled by randomly setting cells per user-specified probabilities during initialization, unlike temporal variations.
 Temporal Drift: Time-dependent conductance changes follow:
@@
 The simulation treats arrays as frozen at the specified time point $t$ for inference calculations.
-In circuit expert mode, statistical models are derived from Monte Carlo circuit analysis or silicon measurements. For SPICE-based characterization, users should:
-  • Create a complete array model including peripheral circuits in a circuit simulator.
-  • Run Monte Carlo simulations incorporating device variations (threshold voltage, device sizing), circuit-level effects (thermal noise, parasitic), and ADC quantization.
-  • Generate statistical distributions by collecting multiple simulation results for each possible MAC output value.
-For silicon-based characterization:
-  1. 1.
-    Collect repeated measurements across multiple array columns.
-  2. 2.
-    Compare expected vs. actual outputs over thousands of trials.
-  3. 3.
-    Build output distributions capturing combined effects of all noise sources.
-The resulting models provide mean and standard deviation pairs for each possible MAC output value.
-The data can be displayed in a ”Confusion Matrix” like those shown in Fig.
-3.
-Users provide these parameters as comma-separated tuples, with one variation value per output level.
-Users store these in the file ’output_noise.csv’.
-During behavioral simulation, the critical loop (Eq.
-3 can be skipped, as the circuit behavior is modeled by the statistical output distribution.
-Therefore, we can calculate ideal partial sums with negligible runtime overhead.
-After ideal partial sums are calculated, the framework samples from the MAC distributions to efficiently inject noise into the outputs.
-This statistical approach enables rapid evaluation and SPICE-level accuracy through empirically validated noise characteristics (see case studies in Section IV.D).
-TABLE I: Design Space Exploration Parameters
-Parameter Range / Values
+Circuit expert mode uses statistical models from Monte Carlo SPICE (varying device properties, temperature, voltage) or silicon measurements (repeated column measurements).
+Both produce mean/standard deviation pairs for each MAC output, visualized as confusion matrices (Fig.
+3) and stored in ’output_noise.csv’.
+This approach skips the critical loop (Eq.
+3), calculating ideal partial sums with minimal overhead, then sampling MAC distributions for noise injection.
+Enables rapid evaluation with empirically validated characteristics (Section IV-B).
+While this statistical approach captures the aggregate impact of these effects, it doesn’t model detailed physics per input/weight.
+This results in a trade-off in physics-accurate simulations for fast large-scale design exploration.
+For highly detailed analysis of such effects, tools like CrossSim [30] may be more suitable.
+Additionally, while the current implementation samples device variations with spatial-independence, users can model spatially-correlated defects by supplying the desired variations for those areas.
+This would require supplying the variation data explicitly for the memory cells in the area of interest along with slight modification to the error injection.
+Such a feature would be useful for post-silicon measured defects and may be explicitly supported in the future.
+III-D Performance Analysis Framework
+NeuroSim V1.5 uses a trace-based interface for PPA estimation. The behavioral simulator saves quantized inputs/weights as CSV traces (Fig. 1) capturing hardware data patterns.
+The C++ PPA estimator auto-generates accelerator floorplans and creates transistor-level circuit models (arrays, ADCs, interconnects, peripherals) calibrated with SPICE and tapeout data [7].
+Using traces, it calculates per-array energy with explicit bit-serial modeling.
+Weights convert to conductance/capacitance states per configuration parameters.
+In NeuroSim V1.5, we extended the hardware analyzer to support charge-based computation with non-volatile capacitive memories.
+This required the addition of new circuit models for nvCap arrays, charge-to-voltage conversion using an amplifier [59], and control circuits for the unique charge-based computation.
+III-E Support for Vision Transformers
+Vision transformer (ViT) architectures are supported in NeuroSim V1.5 using a hybrid ACIM/DCIM approach described in Fig.
+4.
+Linear layers (Q, K, V projections, MLPs) use ACIM arrays in weight-stationary mode.
+Attention score computation ($QK^{T}$) and weighted aggregation use DCIM arrays, as these operations require writing intermediate attention scores as operands for subsequent multiplication—incompatible with NVM write latency and endurance constraints.
+The hybrid architecture maintains NeuroSim’s existing tiled hierarchy.
+Entire tiles are dedicated to either ACIM or DCIM, with identical bit-serial input processing.
+Multi-head attention maps naturally to this structure: each head has independent projection weights, enabling parallel execution across tiles.
+Different layers map to different tiles, and layer-level pipelining allows all tiles to operate simultaneously on different images, maximizing hardware utilization.
+Outputs from each tile are written to the global buffer before being sent to the next layer.
+The global buffer is sized to fit the appropriate activation data from each tile operating in parallel.
+Softmax and GELU are implemented via 8-bit lookup tables (LUTs), storing input-output relationships for the quantized operators.
+The DCIM array architecture follows [3] with circuit modeling from [54]. NeuroSim V1.5 supports both noise-integrated accuracy simulation and PPA evaluation for ViT models.
+Figure 4: NeuroSim V1.5 ViT dataflow.
+Linear projections execute on ACIM tiles; attention score and aggregation use DCIM tiles.
+Numbered circles indicate the operation sequence within one transformer block.
+Layer-level pipelining maps different layers to different tiles, enabling all tiles to execute in parallel on different images.
+TABLE I: Design Space Exploration Parameters and Device Non-idealities Case Study Parameters
+Parameter Range / Values for DSE Values for Noise Case Studies
 Technology Node 22 nm
@@
 […]
-Network / Dataset ResNet-18 / CIFAR-100
-Array Dimensions $32\times 32$ $256\times 256$
-Input / Weight Precision 6b, 8b
 Input Processing bit-serial
-Memory Cell Precision 1b -- 4b
-ADC Precision 3b -- 11b
-  • * The RRAM parameters are configured based on Intel’s 22nm FinFET platform [6]
-III-D Performance Analysis Framework
-To enable accurate power and performance estimation, NeuroSim V1.5 implements a trace-based interface between its behavioral simulation and hardware evaluation components.
-During inference simulation in PyTorch, the framework saves quantized input and weight values as decimal traces in CSV format (shown in Fig.
-1.
-These traces capture the actual data patterns that would flow through the hardware, accounting for all quantization effects and bit slicing operations.
-The hardware estimator (written in C++) automatically generates a floorplan for the entire accelerator based on the network architecture and user-specified parameters.
-It creates detailed transistor-level circuit models for all components (memory arrays, ADCs, interconnects, peripherals, accumulators, etc.), calibrated using SPICE simulations and data from previous tapeouts [7].
-These models enable accurate hardware performance estimation at advanced technology nodes.
-Using traces from the software component, the estimator calculates detailed energy consumption for each array.
-The bit-serial input scheme is explicitly modeled - each input value is processed one bit at a time, with the framework tracking energy consumption for each operation.
-Weight values from the traces are converted to corresponding conductance or capacitance states based on parameters specified in the hardware configuration file.
-In NeuroSim V1.5, we extended the hardware analyzer to support charge-based computation with non-volatile capacitive memories.
-This required the addition of new circuit models for nvCap arrays, charge-to-voltage conversion using an amplifier [61], and control circuits for the unique charge-based computation.
-The hardware estimator uses these models along with device parameters specified in the configuration file to accurately evaluate PPA metrics for nvCap-based accelerators while maintaining compatibility with the existing trace-based interface between behavioral simulation and hardware analysis.
-III-E Evaluation Setup
+Memory Cell Precision 1b - 4b
+Network / Dataset ResNet-18 / CIFAR-100 VGG8 / CIFAR-10 ResNet-18 / CIFAR-100 ResNet-50 / ImageNet 1k Swin-T / ImageNet 1k
+Array Dimensions $32\times 32$ - $256\times 256$ $128\times 128$
+Rows Active all 8 / 32 / 128
+Input / Weight Precision 6b, 8b 8b
+ADC Precision** 3b - 11b full precision
+  • * RRAM parameters based on Intel’s 22 nm FinFET platform [6]
+  • ** ADC full precision is calculated using equations (6) and (7)
+III-F Evaluation Methodology
 Our experimental evaluation consists of three main components: design space exploration of hardware configurations, inference accuracy analysis under various device non-idealities, and runtime performance comparison across open-source platforms.
-The DNN algorithms and proposed NeuroSim framework are implemented using PyTorch, with all simulations performed on a single NVIDIA RTX A6000 GPU with 48GB GDDR6.
-III-E1 Design Space Exploration
+The DNN algorithms and proposed NeuroSim framework are implemented using PyTorch, with all simulations performed on a single NVIDIA RTX A6000 GPU with 48GB GDDR6 memory.
+III-F1 Design Space Exploration
 For the design space exploration, we examine the trade-offs between hardware efficiency and inference accuracy across a range of architectural parameters.
@@
 For a given array configuration, the maximum output value can be calculated as:
-$out_{max}=R(2^{P_{DAC}}-1)(2^{b_{c}}-1)$ (6)
+$out_{max}=R(2^{P_{DAC}}-1)(2^{b_{cell}}-1)$ (6)
 The minimum required ADC precision ($P_{ADC}$) to capture this full range without quantization loss is:
 $P_{ADC}=\lceil log_{2}(out_{max})\rceil$ (7)
-In our design space exploration, we use binary inputs (1b DAC) and we evaluate two ADC precision settings for each configuration:
+In our design space exploration, we evaluate two ADC precision settings for each configuration:
   1. 1.
@@
 Design parameters were swept using automated scripts to explore the trade-offs in accuracy and performance considering input/weight/ADC quantization and array size.
-When reducing ADC precision below the ideal precision, we adopt a clipping approach, where the sensing margins for each analog output state remain the same regardless of ADC precision, and outputs larger than the precision of the ADC are clipped to the maximum ADC output based on its precision.
-We find that this method results in comparable accuracy to other, non-linear quantization methods while being the most practical to implement in hardware.
-It is important to note that higher precision ADCs require smaller sensing margins between levels, which in practice could lead to increased susceptibility to circuit noise and higher error rates.
-The design-space analysis focuses solely on quantization effects, as detailed circuit-level noise characterization data is not available for the multi-bit memory configurations under study.
-For a complete understanding of system-level reliability, future work should incorporate both quantization effects and circuit-level noise measurements.
-III-E2 Device Noise Case Studies
+Results are shown in Fig.
+5 and discussed further in section IV.
+When reducing ADC precision below the ideal precision, we adopt a clipping approach where the sensing margins for each analog output state remain the same regardless of ADC precision.
+Any outputs larger than the precision of the ADC are clipped to the maximum ADC output based on its precision.
+We find that this method results in comparable accuracy to dynamic quantization methods while also being the most practical to implement in hardware.
+III-F2 Device Noise Case Studies
 For analyzing inference accuracy under device non-idealities, we standardize the hardware configuration to isolate the effects of various noise sources and variations.
-Table II presents the fixed parameters used across all device non-ideality studies.
-TABLE II: Device Non-idealities Case Study Parameters
-Parameter Value
-Technology Node 22nm
-Device RRAM
-[…]
-Networks / Datasets VGG8 / CIFAR-10
-ResNet-18 / CIFAR-100
-ResNet-50 / ImageNet 1k
-Swin-T / ImageNet 1k
-Array Dimensions $128\times 128$
-Activated rows 128
-Input / Weight Precision 8b / 8b
-Input Processing bit-serial
-Memory Cell Precision 1b - 4b
-ADC Precision* full precision
-Device Noise Sources** D2D, SAF, state drift
-  • * ADC precision is calculated using equations (6) and (7)
+Table I presents the fixed parameters used across all device non-ideality studies.
 In these case studies, we analyze different sources of device variation including device-to-device variation, stuck-at-faults, and state drift across multiple DNNs and datasets to demonstrate support for multiple networks and noise types.
-III-E3 Circuit Noise Case Studies
+III-F3 Circuit Noise Case Studies
 Next, we demonstrate the power of our improved noise modeling by using circuit noise data characterized both from SPICE characterization and tapeout measurements.
@@
 Mean and standard deviation for each output state are supplied in a comma separated format (.csv file), where each row of the file corresponds to the mean for each given output (post-ADC) and the standard deviation of the output.
-III-E4 Runtime Evaluation
+III-F4 Runtime Evaluation
 To characterize the improvements in runtime over previous versions of NeuroSim, we record the runtime of the inference accuracy portion under several key conditions.
-The main parameters that influence the runtime are the size of the network and dataset used the MLC level, precision of the DAC if one is used, and the input and weight precision.
-IV Results and Case Studies
-IV-A Design Space Exploration
-We demonstrate NeuroSim V1.5’s capability for comprehensive design space exploration using a 22nm RRAM [6] platform as our baseline technology.
-Following the parameters outlined in Table I, we evaluate the fundamental tradeoffs between inference accuracy and hardware efficiency across architectural choices including array size, precision, and data representation.
-Our analysis focuses on architectural optimization without device/circuit noise effects, as these effects are technology-specific and addressed separately in subsequent noise modeling studies.
-TABLE III: Default 22nm RRAM PPA (ResNet18 / CIFAR100)
+These include the network size, input resolution, the MLC level, precision of the DAC (if one is used), and the input and weight precision.
+TABLE II: Default 22 nm RRAM PPA (ResNet-18 / CIFAR100)
 TOPS TOPS/mm2 TOPS/W FPS Accuracy
@@
   • Example performance using 22 nm RRAM [6] and default settings in NeuroSim: $128\times 128$ array size, 7-bit ADC, 8b/8b input/weight precision.
+IV Results and Case Studies
+IV-A Design Space Exploration
+We demonstrate NeuroSim V1.5’s design space exploration capability using a 22 nm RRAM [6] platform.
+Following Table I, we evaluate accuracy-efficiency tradeoffs across architectural choices including array size, precision, and data representation.
 Figure 5: Example design space exploration for 22 nm RRAM.
@@
 Designs with highest TOPS/W use array sizes of $32\times 32$ and $64\times 64$.
-We provide example performance metrics for the default configuration in NeuroSim using the selected RRAM device.
-As shown in Table III, we achieve 11.6 TOPS at 21.3 TOPS/W while processing 7,770 frames per second for CIFAR-100 images.
-Through systematic exploration of the design space, we can find designs that significantly outperform this baseline.
-Results are shown in Fig.
-4.
-From this exploration, we identify several key trends that characterize optimal CIM architectures:
+Table II shows example performance metrics for the default configuration: 11.6 TOPS at 21.3 TOPS/W while processing 7,770 frames per second for CIFAR-100 images.
+Systematic exploration (Fig.
+5) reveals designs that significantly outperform this baseline, identifying key trends for optimal CIM architectures:
   1. 1.
-    ADC Precision: The Pareto-optimal designs cluster in the 5--8 bit ADC precision range.
-Our analysis shows that ADC precision can generally be reduced by 1-bit from the lossless precision without significant accuracy degradation, and by 2-bits in certain configurations.
+    ADC Precision: Pareto-optimal designs cluster in 5-8 bit range.
+ADC precision can be reduced by 1-bit from lossless precision without significant accuracy loss, and by 2-bits in certain configurations.
   2. 2.
-    Array Dimensions: Optimal designs utilize array sizes between $32\times 32$ and $128\times 128$ cells.
-Smaller arrays reduce ADC precision requirements but necessitate additional accumulation circuitry, while larger arrays necessitate larger ADCs, incurring costs in latency and energy.
-The designs that achieve $>$40 TOPS/W used array sizes of either $32\times 32$ or $64\times 64$.
+    Array Dimensions: Optimal designs use array sizes between $32\times 32$ and $128\times 128$ cells.
+Smaller arrays reduce ADC precision requirements but need more accumulation circuitry, while larger arrays require larger ADCs, incurring latency and energy costs.
+Designs achieving >>40 TOPS/W used $32\times 32$ or $64\times 64$ arrays.
   3. 3.
-    Memory Cell Precision: 2-bit and 3-bit MLC configurations dominate the Pareto frontier, offering an effective balance between storage density and programming reliability.
-These configurations achieve the highest energy efficiency (TOPS/W) for any given accuracy target.
-The interplay between these parameters creates distinct architectural sweet spots.
-For example, the 8b/8b/64/6/2 configuration (input/weight/array size/ADC/MLC) achieves 75% accuracy while maintaining high energy efficiency.
-Designs with reduced precision (6b/6b) can push efficiency above 40 TOPS/W at the cost of $\sim 5\%$ accuracy degradation.
-The ability to rapidly evaluate these configurations through NeuroSim V1.5 enables the optimization of architectural parameters based on specific device and circuit designs.
+    Memory Cell Precision: 2-bit and 3-bit MLC configurations dominate the Pareto frontier, balancing storage density and programming reliability while achieving highest energy-efficiency for any given accuracy target.
+Across the explored configurations, Pareto-optimal designs converge toward 6-7 bit ADC precision regardless of other parameter choices.
+This range represents the balance point between ADC overhead (energy, latency, area) and throughput.
+Multiple parameter combinations achieve similar efficiency at this ADC precision: 1-bit cells with 128 rows, 2-bit cells with 64 rows, or 4-bit cells with 32 rows.
+The choice between these configurations then shifts to device multi-state capability and rate of variations.
+Current ADC designs dominate the energy and latency budget in ACIM arrays.
+If more efficient ADC architectures become available—through techniques such as time-domain sensing, or quantization schemes for reduced-precision —the optimal design point would shift toward larger array sizes or higher-precision MLC devices that are currently constrained by ADC overhead.
 Next, we demonstrate the improved noise-modeling capabilities in the new behavioral simulator.
+Figure 6: Inference accuracy of VGG8 on CIFAR-10, ResNet-18 on CIFAR-100, ResNet-50, and Swin-T on ImageNet with different degrees of 1b RRAM device variation.
+”XRow” describes the number of rows activated in parallel.
 IV-B Noise Modeling Case Studies
-Figure 5: Inference accuracy of VGG8 on CIFAR-10, ResNet18 on CIFAR-100, and ResNet50 on ImageNet using V1.5 across different RRAM (single-level cell) variations.
+We evaluated inference accuracy across multiple neural networks using NeuroSim V1.5’s noise modeling capabilities, including D2D variation, memory state drifting, stuck-at-faults and circuit-level MAC output noise.
+We selected network architectures of increasing complexity: VGG8 for CIFAR-10, ResNet-18 for CIFAR-100, ResNet-50 for ImageNet, and Swin Transformer Tiny for ImageNet.
+Following established methodology [61], we excluded the first convolutional layer from quantization.
 IV-B1 D2D Variation
-We evaluated inference accuracy across multiple neural networks using NeuroSim V1.5’s comprehensive noise modeling capabilities, including D2D variation, memory state drifting, stuck-at-faults and circuit-level MAC output noise.
-To systematically assess accuracy degradation, we selected network architectures of increasing complexity: VGG8 for CIFAR-10, ResNet18 for CIFAR-100, ResNet50 for ImageNet, and Swin Transformer Tiny for ImageNet.
-Following established methodology [62], we excluded the first convolutional layer from quantization in all simulations.
-Figure 5 reveals that accuracy degradation becomes more severe with increasing network complexity and dataset difficulty.
-VGG8 on CIFAR-10 maintains reasonable accuracy until LRS variation reaches 20%.
-In contrast, ResNet18 on CIFAR-100 and ResNet50 on ImageNet show significant accuracy drops at just 10% and 5% LRS variation, respectively.
-Most notably, Swin Transformer Tiny demonstrates the lowest tolerance to device-level variations, highlighting the need for additional research into improving transformer model robustness for analog computing implementations.
-Figure 6: Inference accuracy of VGG8 on CIFAR-10 regarding random device conductance drifting, drifting to Gmax, and drifting to Gmin with 1b cells.
+To explore the effect of memory state variation on inference accuracy, we use a 1-bit RRAM baseline design and select several variation rates in the HRS and LRS ranging from 1% to 40%.
+Based on previous measurement results, we use a higher variation in the HRS than the LRS.
+[…]
+Figure 7: Inference accuracy of VGG8/CIFAR-10 regarding random device conductance drifting, drifting to Gmax, and drifting to Gmin with 1b cells.
 IV-B2 Conductance Drift
@@
 7 examines how three different types of conductance drift (modeled by Equation 5) affect inference accuracy.
-First, random drift shows better accuracy retention compared to drifting toward minimum or maximum conductance.
-This improved resilience occurs because cells at minimum conductance cannot drift lower, while cells at maximum conductance cannot drift higher, effectively limiting the total conductance change possible under random drift.
-Second, drifting toward maximum conductance maintains better accuracy than drifting toward minimum conductance.
-This asymmetry arises from Equation 5, where the drift coefficient v determines the drift direction.
-For the same retention time t and drifting speed, drifting toward minimum conductance results in a larger conductance difference compared to drifting toward maximum, leading to a larger deviation of MAC outputs.
-Figure 7: Inference accuracy of VGG8 and ResNet18 under different stuck-at-fault rates.
+Drifting toward maximum conductance (Gmax) maintains the best accuracy retention.
+Random drifting shows the next best resilience; this occurs because cells at minimum conductance cannot drift lower, while cells at maximum conductance cannot drift higher, effectively limiting the total conductance change.
+Drifting toward minimum conductance (Gmin) results in the largest accuracy degradation.
+This asymmetry between Gmax and Gmin drift arises from Equation 5; for the same retention time $t$ and drifting speed $|v|$, drifting toward Gmin can result in a larger absolute conductance change (if starting from a higher state) compared to drifting toward Gmax (if starting from a lower state), leading to a larger deviation of MAC outputs.
 IV-B3 Stuck-at-Faults
-NeuroSim V1.5 adds modeling capabilities for stuck-at-fault (SAF) errors, extending beyond the D2D variation and drift modeling present in V1.4.
+NeuroSim V1.5 adds stuck-at-fault (SAF) error modeling, extending beyond V1.4’s D2D variation and drift modeling.
 Fig.
-7 demonstrates how SAF errors impact inference accuracy across different networks, with SAF rates bounded by realistic RRAM fabrication constraints: 1.75% for LRS and 9.0% for HRS [42].
-Our analysis reveals several key findings.
-First, both 1-bit and 2-bit cells show similar accuracy degradation at equivalent SAF rates, suggesting that cell precision does not significantly influence SAF sensitivity.
-However, network complexity strongly affects SAF tolerance - ResNet-50 on ImageNet loses accuracy even at minimal SAF rates, while VGG8 and ResNet18 maintain limited functionality only at the lowest SAF rates.
-Notably, SAF errors cause more severe accuracy degradation compared to other noise sources like D2D variation or conductance drift.
-This heightened sensitivity to SAF errors aligns with previous research findings [42], highlighting the critical importance of minimizing stuck faults during device fabrication.
-Figure 8: (a) Actual MAC output with 2b FeFET current domain CIM [43] under 50mV Vth variation; (b) Actual MAC output with 2b FeFET charge domain CIM [43] under 50mV Vth variation; (c) Actual MAC output of RRAM CIM tape-out [41]; (d) Actual MAC output with 1b nvCap charge domain CIM [33, 18] under 50mV Vth variation; (e) Inference accuracy of various neural network algorithms across CIM macro A-D.
+8 shows SAF error impact on inference accuracy across different networks, with SAF rates bounded by realistic RRAM fabrication constraints: 1.75% for LRS and 9.0% for HRS [62].
+Both 1-bit and 2-bit cells show similar accuracy degradation at equivalent SAF rates, suggesting cell precision does not significantly influence SAF sensitivity.
+However, network complexity strongly affects SAF tolerance - ResNet-50 on ImageNet loses accuracy even at minimal SAF rates, while VGG8 and ResNet-18 maintain limited functionality only at the lowest SAF rates.
+SAF errors cause more severe accuracy degradation than other noise sources like D2D variation or conductance drift.
+This aligns with previous findings [62], highlighting the critical importance of minimizing stuck faults during fabrication.
+Figure 8: Inference accuracy of VGG8 and ResNet-18 under different stuck-at-fault rates (1b RRAM, 128x128 array, 7b ADC).
 IV-B4 MAC Output Noise
@@
 To demonstrate NeuroSim V1.5’s flexible noise modeling capabilities, we evaluated inference accuracy across four different CIM macros, incorporating both SPICE-characterized circuit data and silicon measurements from fabricated chips.
-CIM A and B [43] employ 2b FeFET technology with different computing modes: CIM A performs current-mode computation with FeFETs directly, while CIM B implements charge-based computation using FeFET-modulated capacitors.
+CIM A and B [34] employ 2b FeFET technology with different computing modes: CIM A performs current-mode computation with FeFETs, while CIM B implements charge-based computation using FeFET-modulated capacitors.
 Their noise characteristics were derived from Monte-Carlo SPICE simulations.
-CIM C [7] provides real silicon validation through direct measurements from a fabricated RRAM-based CIM chip.
+CIM C [7] uses silicon measurements from a fabricated RRAM-based CIM chip.
 For these three designs, each MAC output level is characterized by unique mean and standard deviation values.
@@
 Figs.
-8(a-d) illustrate the statistical distribution of actual versus ideal MAC outputs for representative cases.
+9(a-d) illustrate the statistical distribution of actual versus ideal MAC outputs for representative cases.
 While CIM A, B, and D show tighter error bounds compared to CIM C, all designs achieve similar accuracy on smaller networks.
 Fig.
-8(e) compares inference accuracy across architectures of increasing complexity.
-VGG8 and ResNet18 maintain reasonable accuracy across all implementations, with CIM C showing slightly higher degradation ( 2%) due to increased output variation.
-For ResNet50, only CIM B and D - which achieve lower standard deviations in their MAC errors - maintain high accuracy.
+9(e) compares inference accuracy across architectures of increasing complexity.
+VGG8 and ResNet-18 maintain reasonable accuracy across all implementations, with CIM C showing slightly higher degradation ( 2%) due to increased output variation.
+For ResNet-50, only CIM B and D - which achieve lower standard deviations in their MAC errors - maintain high accuracy.
 The Swin Transformer exhibits particular sensitivity to output noise, with only the nvCap-based CIM D achieving software-comparable performance.
-These results validate NeuroSim V1.5’s noise modeling framework as a versatile platform for evaluating CIM designs using either simulation data or silicon measurements.
+These results also validate the circuit expert mode’s statistical modeling approach.
+By characterizing MAC output distributions from SPICE simulations or silicon measurements and injecting noise accordingly, the framework captures the aggregate impact of device and circuit non-idealities without requiring per-cell physical simulation.
 By providing a standardized methodology for capturing MAC output statistics, the framework enables direct comparison of different technologies and compute circuits under realistic operating conditions.
-IV-C Runtime Evaluation and Simulator Comparison
-We first compare NeuroSim V1.5 against existing open-source ACIM simulation frameworks to highlight key capabilities and performance improvements (Table IV).
-CrossSim [37] provides detailed device-level accuracy simulation but faces significant runtime overhead when modeling noise effects.
-AIHWKit [63] specializes in noise-aware training but offers limited support for general noise modeling and architectural exploration.
-While NeuroSim V1.4 [54] includes PPA analysis and has some noise modeling, it was constrained by slower inference speeds and limited network support.
+Figure 9: (a) Actual MAC output with 2b FeFET current domain CIM [34] under 50mV Vth variation; (b) Actual MAC output with 2b FeFET charge domain CIM [34] under 50mV Vth variation; (c) Actual MAC output of an RRAM tape-out [33]; (d) Actual MAC output with 1b nvCap charge domain CIM [27, 18] under 50mV Vth variation; (e) Inference accuracy of various neural network algorithms across CIM macro A-D.
+IV-C Noise Sensitivity Analysis of ViTs
+As shown in Fig.
+6, Swin-T exhibits lower tolerance to analog noise compared to ResNet-50.
+To understand the underlying mechanisms, we analyze errors at both the layer and circuit levels using device-to-device (D2D) variations as the primary noise source.
+IV-C1 Related Work on Transformer Sensitivity
+Prior work attributes transformer sensitivity in analog CIM to non-uniform activation distributions causing quantization challenges [50, 49], heightened input/output interface noise susceptibility [63], and MSB-cycle error amplification in bit-serial computation [64].
+IV-C2 Layer-wise Error Analysis
+To identify whether specific layers or architectural components drive Swin-T’s higher noise sensitivity, we compute the root mean square error (RMSE) in each layer output under the 4%/2% D2D variation condition.
+RMSE is computed as $\sqrt{\text{mean}((\hat{y}-y)^{2})}/\sqrt{\text{mean}(y^{2})}$, where $y$ is the ideal floating-point layer output and $\hat{y}$ is the noisy output.
+Tests were conducted for both ResNet-50 and Swin-T on ImageNet with 1-bit cells, an array size of $\times128$, and 7-bit ADC.
+Figure 10: Layer-wise RMSE for ResNet-50 vs Swin-T under 4%/2% (HRS/LRS) D2D variation.
+Swin-T shows higher error variance with attention layers (triangles) exhibiting greater sensitivity than non-attention layers (squares).
+[…]
+Figure 11: CIM ADC output distribution under ResNet-50 and Swin-T using 8b input and 8b weight with 128x128 arrays and 4%/2% (HRS/LRS) D2D variations. Collected from 100 images in ImageNet dataset.
+IV-C3 ADC Error Analysis
+The layer-level RMSE analysis reveals error magnitudes after accumulating partial sums and dequantizing to floating-point.
+To understand the root cause of these errors, we examine errors at the ADC output stage, where integer MAC partial sums are quantized.
+We define error rate as the percentage of partial sums where the noisy ADC output deviates from the ideal value, computed per expected output bin.
+Analysis of ADC output distributions reveals that Swin-T generates significantly higher average ADC outputs compared to ResNet-50 (Fig.
+11).
+Larger ADC outputs experience increased error rates due to D2D variations, as shown in Fig.
+12.
+Consequently, Swin-T accumulates a higher fraction of erroneous ADC outputs across all layers, contributing significantly to accuracy degradation.
+The increased error rate at higher ADC output values stems from the accumulation of device variations across activated cells.
+Each memory cell contributes independent D2D variation to the column current/charge.
+When more cells are activated (corresponding to larger partial sums), the total variance increases with the number of contributing devices.
+Swin-T’s activation and weight distributions produce systematically higher partial sums than ResNet-50, placing a larger fraction of computations in the high-error region of Fig.
+12.
+Prior work has attributed transformer sensitivity in analog CIM primarily to MSB error amplification during bit-serial computation [64].
+While MSB errors do have a disproportionate impact on output magnitude, this mechanism alone does not explain why ViTs exhibit lower noise tolerance than CNNs—both architectures use identical bit-serial processing.
+Our analysis indicates that the distribution of ADC outputs, determined by activation sparsity and weight magnitude statistics, is the distinguishing factor.
+CNNs with ReLU activations produce sparser outputs concentrated at lower ADC values, while transformers with GELU activations and larger weight magnitudes shift the distribution toward higher values where error rates increase.
+Figure 12: Error rate as a function of expected ADC output value for ResNet-50 and Swin-T with 128x128 arrays under 4%/2% (HRS/LRS) D2D variations. Collected from 100 images on the ImageNet dataset.
+IV-C4 Mitigation Strategies
+To improve Swin-T accuracy under analog noise, we decreased the size of the memory arrays to $32\times 128$.
+This effectively reduces the ADC outputs to values less than 32, where the rate of errors is below 2% (Fig.
+12).
+As shown in Fig.
+6, this modification achieves accuracy close to the software baseline at 4%/2% D2D variation; however, accuracy quickly falls off at higher variation.
+To achieve accuracy comparable to ResNet-50 under D2D variations, we need to reduce the number of active rows in parallel to 8.
+The reduction in active rows does not require a new hardware configuration, as each group of 8 rows in an array are activated sequentially.
+We quantify the tradeoff in performance in (Table III).
+We find that Swin-T using $32\times 128$ configuration achieves near parity with ResNet-50 in throughput and energy efficiency but with $5.4\times$ worse area efficiency.
+While comparable accuracy is achievable, CNNs remain preferable for CIM-based vision accelerators.
+Alternative approaches to improve transformer resilience to ACIM noise include: (1) hardware-aware training with noise injection [65]; (2) input and weight rescaling methods to shift their distributions to fit within quantization constraints [66]; (3) Training low-rank adapters to correct errors arising from CIM noise [67].
+ResNet-50’s inherent advantages—weight distributions closer to zero and ReLU-induced sparsity—naturally produce lower ADC outputs and improved noise tolerance.
+Applying techniques to encourage these input and weight distributions in transformer architectures can bridge this robustness gap without requiring hardware modifications.
+TABLE III: 22 nm RRAM ResNet-50 and Swin-T PPA under different row parallelism
+Model Array Read Rows TOPS TOPS / mm2 TOPS/W FPS
+ResNet50 128x128 128 4.84 0.0147 5.04 378
+Swin-T 32x128 32 4.83 0.00272 4.84 388
+Swin-T 32x128 8 3.46 0.00192 4.34 278
+IV-D PPA Breakdown of Swin-T
+Next, we evaluate the PPA of Swin Transformer [68] on ImageNet.
+The linear and convolution layers are implemented in 22 nm RRAM-based ACIM while the Q,K,V matrix multiplications in attention are computed in 22 nm SRAM-based DCIM.
+Both CIM arrays are configured with a size of $32\times 128$.
+Fig.
+13 presents the detailed PPA breakdown.
+For area, DCIM accumulation including adder trees and shift adders dominates, resulting in $1.5\times$ larger area than ACIM.
+However, the energy profile differs: the ADC in ACIM contributes the largest portion of energy consumption compared to significantly lower consumption of DCIM computation.
+This is because the linear and convolutional layers have higher arithmetic intensity than attention, requiring a higher usage of the ACIM arrays and thus higher energy consumption.
+This asymmetry suggests that for area-constrained designs, minimizing DCIM-computed operations (e.g., offloading some attention to external processors) may be more effective, while for energy-constrained designs, ADC optimization or reduced-precision sensing offers larger gains if accuracy can be maintained via quantization-aware tuning methods.
+For the critical path latency, the adder tree in DCIM and the ADC in ACIM show comparable delays.
+The peripheral circuits in ACIM array introduce additional overhead with longer overall latency compared to DCIM.
+A potential optimization for DCIM-based attention is leveraging finer-grained pipelining at the window level or head level with carefully designed dataflow and higher memory array reuse, which can reduce area while maintaining high-throughput.
+Figure 13: PPA breakdown of Swin-T (25M parameters).
+The majority of area is consumed by DCIM adder trees within memory arrays.
+Array size was reduced to 32x128 to decrease adder tree size.
+The majority of energy consumption and latency arises from ACIM ADC.
+IV-E Runtime Evaluation and Simulator Comparison
+Finally, we compare NeuroSim V1.5 against existing open-source ACIM simulators (Table IV).
+CrossSim [30] excels at device-level physical simulation with high fidelity but higher runtime, especially for complex noise.
+AIHWKit [63] focuses on noise-aware training with phase-change memory.
+NeuroSim V1.4 [54] introduced integrated PPA and some noise modeling but had slower inference and narrower network support.
+NeuroSim V1.5 distinguishes itself by integrating flexible, gpu-accelerated noise modeling, and vision transformer support, with circuit-level PPA estimation.
+This enables a wider range of noise, quantization, and circuit parameters to be explored than in other open-source simulators.
 TABLE IV: Comparison of DNN ACIM Simulators
@@
 nvCap
-Supported network types*
+Supported Network Types
 Linear,
@@
 Transformer
-Custom Compute Circuit Support No No No Yes
 Bit Slicing Support Yes No Yes Yes
 PPA Support No No Yes Yes
-  • * PPA evaluation excluded.
-The runtime is primarily influenced by the precision parameters that determine computation complexity through Equation 3: DAC precision affects the number of input cycles, cell bit-precision (MLC) determines the number of weight components, and noise modeling adds statistical sampling overhead.
-With 8b/8b/None configuration (DAC / MLC / Noise) shown in Table V, V1.5 achieves a baseline runtime of 0.95 ms/image for VGG8 and CIFAR-10.
-When using 1b cells to represent 8b weights, runtime increases $\sim$4$\times$ to 4.1 ms/image due to additional bit-slicing operations.
-Adding device noise results in negligible runtime increase and adding per-MAC output noise increases runtime to only 5.1 ms/image.
-NeuroSim V1.5 achieves this dramatic improvement in simulation efficiency through three key advances: a PyTorch-optimized ACIM behavioral simulator, noise modeling with pre-characterized statistical noise modeling and TensorRT integration for efficient quantization.
-As shown in Fig.
-9, these optimizations deliver a 6.51$\times$ speedup over V1.4 for CIFAR-10 inference with 1-bit cells.
+Table V details runtime scaling with precision and noise configuration.
+NeuroSim V1.5 achieves 0.95 ms/image baseline for VGG8/CIFAR-10, increasing to 4.1 ms/image with 1b cells due to bit-slicing overhead.
+Device noise adds negligible cost; per-MAC output noise increases runtime to 5.1 ms/image.
+[…]
 Figure 14: Inference runtime of different neural network algorithms using V1.4 and V1.5 across 1b/2b/4b RRAM cell states without device non-idealities and noise.
-Tables V and VI further demonstrate V1.5’s efficient noise modeling capabilities on larger networks towards CIFAR-100 and ImageNet datasets.
-Compared to CrossSim, which shows a significant slowdown when incorporating device noise, V1.5’s statistical approach adds minimal overhead - just 1.3$\times$ for uniform noise and 3.1$\times$ for output-dependent noise simulation.
-This efficiency stems from our pre-characterized noise distribution approach, eliminating the need for detailed circuit simulation during inference discussed in section III-B.
 TABLE V: Inference Runtime (Seconds/Image) Comparison of VGG8 on CIFAR-10 (8b activation, 8b weight) between NeuroSim V1.4 and NeuroSim V1.5
 DAC Precision/
-/MLC/Noise
+MLC/Noise
 NeuroSim
@@
   • * 0.039: same noise on each MAC output, 0.085: individual noise on each MAC output
-These advances establish NeuroSim V1.5 as a comprehensive development platform that bridges detailed device/circuit-level analysis with system-level architectural exploration.
-The framework’s efficiency and adaptability make it particularly valuable as the field moves toward more complex network architectures requiring extensive accuracy and performance optimization.
-TABLE VI: Inference Runtime (Seconds/Image) Comparison of ResNet50 on ImageNet (8b activation, 8b weight) between CrossSim, AIHWKit, and NeuroSim V1.5
+[…]
+These advances improve NeuroSim’s functionality as a design space exploration platform across memory devices, computing methods, and technology nodes.
+Evaluating both network accuracy and performance is essential for optimal CIM hardware design, as ignoring either metric provides an incomplete picture and can exaggerate throughput and energy-efficiency.
+TABLE VI: Inference Runtime (Seconds/Image) Comparison of ResNet-50 on ImageNet (8b activation, 8b weight) between CrossSim, AIHWKit, and NeuroSim V1.5
 DAC Precision/
-/MLC/Noise
+MLC/Noise
 CrossSim
@@
   • * 0.29: same noise on each MAC output, 0.61: individual noise on each MAC output
-V Conclusion
-The growing complexity of AI models places increasing demands on CIM accelerator design, requiring tools that can efficiently evaluate both accuracy and hardware implementation trade-offs.
-This work presents NeuroSim V1.5, which combines accuracy analysis with power, performance, and area estimation to support systematic design space exploration of CIM architectures.
-The framework’s integration with TensorRT enables evaluation of pre-trained networks without modification, while our flexible noise modeling methodology supports both detailed device-level analysis and efficient MAC-level noise injection, validated against published SPICE simulations and silicon measurements.
-These capabilities, combined with up to 6.5× faster runtime compared to NeuroSim V1.4, enable more extensive design space exploration than previous versions.
-Through our case studies with a variety of device technologies and neural network topologies, we demonstrate how NeuroSim V1.5 can provide insights into CIM design trade-offs:
+V Conclusion and Future Work
+The growing complexity of AI models places increasing demands on CIM accelerator design, requiring tools that can efficiently evaluate both accuracy and hardware performance.
+This work presents NeuroSim V1.5, which combines accuracy analysis with power, performance, and area estimation.
+A key contribution is the heterogeneous ACIM/DCIM architecture extending CIM support to transformer models—increasingly important as transformers dominate AI workloads from vision to language.
+[…]
+Through our case studies with various device technologies and neural network topologies, we demonstrate how NeuroSim V1.5 can provide insights into CIM design trade-offs:
   1. 1.
-    Optimal designs must balance array size, ADC precision, and multi-level cell capability.
+    Pareto-optimal designs converge around the 5-7 bit ADC configuration, underscoring the importance of ADC overhead in ACIM architectures.
   2. 2.
-    Network complexity influences sensitivity to device-level variations, particularly in larger networks like ResNet-50 and Swin-T.
+    Heterogeneous ACIM/DCIM architectures enable transformer acceleration; however, limiting DCIM circuits to ACIM-compatible technology nodes reduces area efficiency.
+Heterogeneous chiplet integration with separate ACIM and DCIM dies could offer a solution.
   3. 3.
-    Charge-domain computation (e.g., based on nvCap) is a promising device for future large-scale ACIM acceleration.
-Looking forward, the emergence of large language models (LLMs) with billions to trillions of parameters presents new challenges beyond on-chip memory capacity.
-These models will require GB-level high bandwidth memory (HBM) and process-in-memory (PIM) architectures that place compute logic near DRAM dies.
-While direct LLM and PIM support is beyond NeuroSim V1.5’s current scope, we have initiated efforts toward these capabilities through our work on 3D-stacked DRAM on TPU-like architectures [64, 65].
+    ViTs exhibit lower noise tolerance and area efficiency than CNNs of comparable size. For vision applications where both architectures are viable, CNNs remain better suited for CIM hardware.
+Looking forward, several enhancements are planned: lookup table-based noise modeling for complex effects like IR drop, generalizing the PPA framework for flexible system-level exploration similar to CIMLoop, and incorporating time-domain input encoding for broader CIM coverage.
+Extending NeuroSim to support large language models is a primary future direction, as transformers are the de-facto architecture for these workloads.
+LLMs present unique challenges requiring GB-level parameter storage—addressable through multi-die CIM integration or process-in-memory (PIM) architectures with CMOS compute near high-bandwidth memory.
+We have initiated PIM efforts through our work on 3D-stacked DRAM architectures [69, 70].